### v1

# AI-Empowered Integrated Sensing and Communications

Abstract
Integrating sensing and communication (ISAC) can help overcome the challenges of limited spectrum and expensive hardware, leading to improved energy and cost efficiency. While full cooperation between sensing and communication can result in significant performance gains, achieving optimal performance requires efficient designs of unified waveforms and beamformers for joint sensing and communication. Sophisticated statistical signal processing and multi-objective optimization techniques are necessary to balance the competing design requirements of joint sensing and communication tasks. Since model-based analytical approaches may be suboptimal or overly complex, deep learning emerges as a powerful tool for developing data-driven signal processing algorithms, particularly when optimal algorithms are unknown or when known algorithms are too complex for real-time implementation. Unified waveform and beamformer design problems for ISAC fall into this category, where fundamental design trade-offs exist between sensing and communication performance metrics, and the underlying models may be inadequate or incomplete. This article explores the application of artificial intelligence (AI) in ISAC designs to enhance efficiency and reduce complexity. We emphasize the integration benefits through AI-driven ISAC designs, prioritizing the development of unified waveforms, constellations, and beamforming strategies for both sensing and communication. To illustrate the practical potential of AI-driven ISAC, we present two case studies on waveform and beamforming design, demonstrating how unsupervised learning and neural network–based optimization can effectively balance performance, complexity, and implementation constraints.
Index Terms:
Integrating sensing and communications (ISAC), AI, beamforming, waveform, constellation, radar, sensing, performance metrics, unsupervised learning, autoencoder, 6G.
I Introduction
With the goal of “everything is sensed, everything is connected, and everything is intelligent”, the sixth-generation (6G) wireless systems will incorporate intelligent sensing to enable innovative services such as localization, autonomous vehicles, area imaging, activity detection, identification, and monitoring [1]. Historically, radar sensing and communication have been designed and operated independently, thus using different frequency bands. Integrated sensing and communications (ISAC) enables spectrum and hardware sharing to reduce the cost and resource demands. According to the federal communications commission (FCC) [2], the coexistence of radar and wireless systems can be observed in the L-band (1-2 GHz), S-band (2-4 GHz), C-band (4-8 GHz), and millimeter-wave (mmWave) band (30-300 GHz).
ISAC exemplifies some of the key distinctions between 5G and 6G. In 6G, communication signals will also serve as a medium for sensing, allowing the network to simultaneously detect, track, and map its environment while transmitting data. For example, ISAC enables autonomous driving and unlocks new use cases such as environmental awareness, precise localization, and context-aware services, pushing beyond the capabilities of 5G, where sensing remains a separate function. In addition, ISAC leads to greater spectral and energy efficiency, reduces congestion, and prevents spectrum under-utilization [3, 4, 5]. However, the true potential of ISAC technology can be realized through efficient designs of unified waveforms and beamformers. Nonetheless, achieving optimal performance is challenging, in part because sensing and communication have contrasting objectives and performance metrics.
The performance of sensing tasks—such as estimation, detection, and recognition—is typically quantified via mean squared error (MSE), Cramér-Rao lower bound (CRLB), and receiver operating characteristic (ROC) [3, 4, 5]. On the other hand, the performance of wireless communications is mainly characterized by spectral and energy efficiency and reliability metrics. Consequently, there are fundamental trade-offs between the performance metrics of sensing and communications. In particular, in a Gaussian channel, the main metric for communication data rates—the mutual information (MI)—is related to the minimum mean square error (MMSE) of sensing by $\frac{\partial}{\partial\gamma}I(\gamma)=0.5\mathrm{MMSE}(\gamma),$ where $\gamma$ represents the signal-to-noise ratio (SNR)[6]. Thus, for a Gaussian channel, Gaussian inputs provide an upper bound on both the mutual information and MMSE. In other words, the Gaussian input is the most favorable for communication (as it maximizes $I(\gamma)$), yet the least favorable for sensing (as it maximizes $\mathrm{MMSE}(\gamma)$). This is demonstrated in Fig. 1, where the $\rm MMSE$ and $\rm MI$ are also plotted for two practical constellations, namely binary phase-shift keying (BPSK) and quadrature phase-shift keying (QPSK). It is clear that QPSK is preferred if mutual information is the metric, whereas BPSK is preferred when MMSE is a concern.
Figure 1: Mutual information ($I$ in nats per channel-use) and MMSE versus SNR for an AWGN channel with three different input signals: Gaussian, QPSK, and BPSK. While Gaussian signaling achieves the maximum information rate, it also results in the highest (worst) MMSE.
Given the fundamental trade-offs and the inherent multi-objective optimization requirements, advanced statistical signal processing and multi-objective optimization techniques are necessary to balance the competing design objectives of joint sensing and communication. Significant research efforts have been dedicated to achieving the goals of ISAC through classical model-based designs and theoretical analyses, leveraging communication and statistical signal processing theories [7, 1, 3, 8, 9, 10]. However, designing optimal unified waveforms and beamformers using model-based analytical approaches remains often prohibitively complex and challenging.
Machine learning (ML), and its subset deep learning (DL), has revolutionized wireless physical layer designs [11]. Deep learning is a powerful tool for extracting channel features and leveraging them in various tasks through universal model approximations and solving non-linear multi-objective optimization problems. The use of DL/ML offers significant advantages in this context by enabling performance optimization without relying on explicit models. Instead, DL approaches leverage implicit learning to extract channel parameters directly from high-dimensional data, making them well-suited for the dynamic and unpredictable conditions inherent to ISAC systems.
I-A Motivation
Driven by the limitations of model-based analytical approaches in designing efficient unified beamforming and waveform solutions, this tutorial aims to establish the theoretical and algorithmic foundations of artificial intelligence (AI)-assisted waveform and beamformer designs. These AI-driven approaches can unify and optimize ISAC functionality while enhancing ISAC integration gains. A critical question is when and why learning-based approaches are needed. We answer this in the following.
I-A1 Imperfect Modeling of Mismatches
In the context of ISAC, learning-based approaches become essential when traditional model-based methods fail to account for imperfections and mismatches in models or hardware. These mismatches can stem from imperfect (e.g., estimated or quantized) channel state information (CSI), dynamic environments, interference, hardware limitations, or the inherent complexity of integrating sensing and communication. Learning-based techniques, such as deep learning or reinforcement learning, can adapt to these uncertainties by leveraging real-time data, enabling robust waveform and beamformer designs. This adaptability is crucial for optimizing ISAC performance in scenarios where precise analytical models are either insufficient or computationally intractable.
I-A2 End-to-End Learning
Deep autoencoder-based end-to-end learning [12] offers a powerful approach to system design, as it enables the joint optimization of multiple components without relying on explicit models, making it highly adaptable to the dynamic and uncertain environments characteristic of ISAC. The broad impact and benefits of DL, and autoencoders in particular, are mainly due to their effectiveness in processing high-dimensional data with embedded low-dimensional structures [13]. This characteristic makes them particularly attractive for ISAC problems such as CSI estimation, interference management, constellation design, and waveform design.
For example, from a wireless communications perspective, received pilot signals represent high-dimensional observations typically acquired through sparse mmWave and sub-terahertz (THz) channels, which inherently possess low-dimensional structure. However, the number of such observations is limited by the channel coherence interval, which is constrained by delay and Doppler spreads. Besides, when hybrid beamforming is employed, the baseband output becomes a nonlinear, low-dimensional transformation of these high-dimensional observations, making accurate estimation of radar and communication channel parameters highly challenging.
I-A3 6G Networks Are Underpinned by AI
6G networks are expected to be AI-driven. Embedded AI into networking equipment enables the network to learn, self-manage, operate autonomously, and reduce costs. AI-driven 6G networks leverage machine learning at every layer—from physical layer optimization to network management—facilitating seamless integration of sensing and communication functionalities. ISAC is anticipated to play a pivotal role in 6G systems [3, 14, 4], advanced vehicle-to-everything systems [15, 16, 17], smart industrial Internet of things (IoT) networks [18, 19, 20, 21], and next-generation wireless local area networks [22, 23]. Applications of ISAC include, but are not limited to, localization, tracking, autonomous vehicles, area imaging, human activity detection, environmental monitoring, and surveillance [3, 24, 10]. Incorporating AI into ISAC further enhances the development of highly adaptive, intelligent, and efficient communication systems.
I-B Scope and Related Tutorials
This work presents a tutorial dedicated to AI-based ISAC, with a primary focus on how AI can enhance ISAC design through the joint optimization of waveform and beamforming. We discuss various ML/DL–based optimization approaches for this joint design, highlighting their advantages over traditional analytical and numerical methods used for multi-objective optimization.
Several papers have surveyed the past research contributions to the advancement of ISAC technology [24, 25, 26, 27, 1, 28, 29, 30, 31, 32, 33, 34]. However, none of those tutorials and surveys are dedicated to AI in ISAC. In [1], a survey on perceptive mobile networks explores ISAC functionalities beyond localization, tracking, and object recognition. Such networks serve as radio sensing platforms, enabling smart sensing applications while ensuring uncompromised communication. In [16, 24], the fundamental performance limits of ISAC have been studied. In particular, [24] categorizes existing techniques as device-based ISAC, device-free ISAC, device-free sensing, and device-based sensing. A survey on orthogonal frequency-division multiplexing (OFDM) and orthogonal time frequency space (OTFS) modulation for ISAC is presented in [25].A more comprehensive survey on ISAC waveform and multiple access design can be found in [33]. In [26], an overview of network sensing and sensing-aided communication techniques for cellular systems is provided. Fundamental limits of ISAC are presented in [34].
The integration gains, system and signal design, network resource management for integrated sensing, communication, and computation are discussed in [27]. A survey of interference management techniques for ISAC is presented in [28]. Multi-functional reconfigurable intelligent surface-aided ISAC systems are reviewed in [29]. In [30], intelligent multi-modal sensing-communication integration is surveyed using the concept of machine synesthesia for dynamic environments. A survey on secure ISAC for intelligent connectivity in 6G networks is presented in [31], emphasizing the need for dynamic adaptability and self-learning capabilities. Finally, [32] provides a recent tutorial on sensing with communication signals.
I-C Contribution
This tutorial provides a comprehensive exploration of ISAC systems, with a particular focus on AI-driven advancements in the context of 6G networks. It emphasizes the limitations of traditional modeling approaches and showcases the potential of various AI-based techniques—such unsupervised learning—to overcome these challenges. The growing role of AI in next-generation networks is discussed, leading to an in-depth examination of AI-based ISAC design.
We begin with a primer on ISAC, introducing key concepts in both communication and sensing, along with performance metrics relevant to each domain. Radar technology is presented as a a core sensing component, and its integration with communication systems is discussed. We then frame ISAC problems as multi-objective optimization tasks, comparing conventional approaches—such as weighted sum, $\epsilon$-constraint, and Pareto optimization—with modern learning-based solutions, including multi-task learning and multi-objective learning.
A major contribution of this tutorial is its detailed examination of waveform and beamforming design for ISAC and the role AI plays in enhancing it. We cover traditional beamforming techniques in communication, radar, and ISAC and explore learning-based approaches, including supervised, unsupervised, and reinforcement learning, as well as model-based learning techniques like plug-and-play methods and algorithm unrolling.
To make AI-driven ISAC designs more tangible, we present two concrete case studies on AI-empowered waveform and beamforming design.
  • The first case study explores unsupervised learning-based waveform design, and demonstrates how the sensing-communication trade-off can be captured via a custom loss function, with constraints enforced using a proper layer. Performance is evaluated and compared to existing approaches using achievable sum rate, probability of detection, and transmit beampatterns, while also analyzing computational complexity and design trade-offs.
  • The second case study presents a hybrid beamforming method using a learned optimizer, which optimizes analog and digital precoders by reformulating the problem as constrained least-squares optimization, and solves it via neural network-based projection. Numerical results show that the proposed unrolled hybrid beamforming method achieves spectral efficiency close to optimal digital beamforming, particularly with more receiving antennas.
These studies underscore AI’s potential for efficient and scalable waveform design and beamforming in next-generation wireless systems.
Finally, we outline future research directions and open problems, such as the development of end-to-end ISAC systems, the application of transformer neural networks for unified waveform and beamformer design, and digital twin-aided ISAC architectures. By bridging the gap between traditional signal processing and modern AI-driven techniques, this tutorial serves as a valuable resource for researchers and practitioners at the forefront of ISAC technology.
TABLE I: List of Key Abbreviations.
Acronym Description
3GPP 3rd Generation Partnership Project
6G Sixth generation
AI Artificial intelligence
AWGN Additive white Gaussian noise
BER Bit error rate
BS Base station
CRLB Cramér-Rao lower bound
CSI Channel state information
CW Continuous wave
DBF Digital beamforming
DFRC Dual-functional radar-communication
DL Deep learning
DNN Deep neural network
EVM Error vector magnitude
FMCW Frequency modulated continuous wave
GNN Graph neural network
HBF Hybrid beamforming
IoT Internet of things
ISAC Integrated sensing and communications
LiDAR Light detection and ranging
LSTM Long short-term memory
MI Mutual information
MIMO Multiple-input multiple-output
ML Machine learning
MMSE Minimum mean squared error
MOL Multi-objective learning
MSE Mean squared error
MTL Multi-task learning
MUI Multi-user interference
OFDM Orthogonal frequency division multiplexing
OTFS Orthogonal time frequency spacing
PMCW Phase modulated continuous wave
QPSK Quadrature phase-shift keying
RF Radio frequency
ROC Receiver operating characteristic
SDR Semi-definite relaxation
SER Symbol error rate
SINR Signal-to-interference-plus-noise ratio
SNR Signal-to-noise ratio
SVD Singular value decomposition
II A Primer to ISAC
In 6G, communication signals will not only facilitate data transmission but will also serve as a medium for sensing and radar applications. This dual functionality will enable next-generation networks to simultaneously detect, track, and map its surrounding environment while maintaining reliable communication. By integrating sensing capabilities directly into the communication infrastructure, 6G networks will support advanced use cases such as environmental monitoring, high-resolution imaging, and object detection. This convergence of communication, radar, and sensing technologies will pave the way to new possibilities for applications like autonomous driving, smart cities, and new IoT systems, where real-time situational awareness and efficient data exchange are critical.
Sensing and radar are closely related concepts, particularly in the field of remote sensing and detection technologies. Below, we explain each of these topics, their applications, their interrelation, and their integration with communication technologies. We also discuss the typical performance metrics used in both communication and radar systems.
II-A Communications
Communication systems involve the transmission of signals, such as voice, video, and data from a sender to a receiver through various physical media, including cables, optical fibers, or wireless channels. These systems rely on a series of processes, such as source and channel encoding, modulation, multiplexing, beamforming and waveform design to prepare the information for efficient transmission over the selected medium. At the receiver’s end, the reverse operations are applied to reconstruct the original signal.
The primary objective of a communication system is to ensure the fast and reliable exchange of information by maximizing the possible data transmission rate while minimizing the possible error rate. Achieving this involves addressing challenges like noise, interference, channel distortions, hardware imperfectness, and synchronization, which can degrade signal quality. Modern communication systems also aim to optimize factors such as bandwidth efficiency, power consumption, and latency to meet the demands of various applications, from high-speed internet and mobile communications to streaming and real-time conferencing.
5G networks are designed to support diverse applications by providing enhanced performance metrics. Building upon the advancements of 5G, 6G technology aims to significantly enhance key performance indicators. Projected targets for 6G include a peak data rate of 1 terabit per second, an experienced data rate of 1 gigabit per second, and a peak spectral efficiency of 60 bits per second per Hertz (bits/s/Hz) [35]. Additionally, 6G aspires to achieve user plane latency as low as 100 microseconds ($\mu$s) and reliability levels reaching $10^{-7}$. These ambitious goals are designed to support applications such as immersive communications, hyper-reliable low-latency services, autonomous driving, and the extensive connectivity required for the IoT.
II-A1 Communications Performance Metrics
The main figure of merit in digital communication is the energy per bit to noise power spectral density ratio, ${E_{b}}/{N_{0}}$, where $E_{b}$ denotes the energy per bit and $N_{0}$ corresponds to the noise power spectral density. It is related to the SNR by ${E_{b}}/{N_{0}}=\text{SNR}\cdot{W}/{R_{b}}$, where $W$ denotes the system bandwidth, and $R_{b}$ is the bit rate [36]. The ratio ${R_{b}}/{W}$ is known as bandwidth efficiency, or spectral efficiency, and is measured in bits/s/Hz. With ideal pulse shaping, for $M$-ary digital modulation schemes like QPSK this efficiency is $\mathrm{log}_{2}M$ [37]. Another important performance metric is the bit error rate (BER), which represents the probability of bit errors in the received data. BER is commonly plotted as a function of either $\frac{E_{b}}{N_{0}}$ or SNR.
The achievable data rate refers to the data transmission rate in a channel, considering practical constraints. The capacity of a channel is defined as the highest achievable data rate when the BER goes to zero. For an additive white Gaussian noise (AWGN), Shannon’s formula $\rm C=W\,\log[2]{1+{SNR}}$ defines the capacity in bits/s. There are other important metrics such as throughput, outage probability, error vector magnitude (EVM), signal-to-interference-plus-noise ratio (SINR), and latency, all defined in Table II.
TABLE II: Common Performance Metrics for Communication Systems
Metric Definition Description & Additional Information
BER Probability of bit errors in received data. Lower BER indicates more reliable communication. BER depends on ${E_{b}}/{N_{0}}$ (or SNR), modulation scheme, and coding.
SNR Ratio of signal power to noise power. Higher SNR leads to better detection and lower BER.
SINR Ratio of signal power to the sum of interference and noise power. Higher SINR results in better communication quality, especially in interference-limited networks.
Spectral efficiency Bits transmitted per second per Hz of bandwidth. Higher spectral efficiency allows for more data transmission in limited bandwidth.
Capacity Theoretical upper bound on data transmission rate. Given by Shannon’s formula $\rm C=W\,\log[2]{1+SNR}$, represents the maximum possible data rate with vanishing errors.
Achievable rate Data transmission rate considering practical constraints. Determined by modulation, coding, and BER requirements. It is lower than capacity because of implementation limitations.
Throughput Actual data successfully transmitted and received. Affected by network congestion, retransmissions, and protocol overhead. It is always lower than the achievable data rate.
Outage probability Probability that the received signal falls below a threshold. Important in fading channels to assess communication reliability.
Frame error rate Probability of errors in transmitted frames. Used in packet-based communication to assess reliability.
Latency Delay in data transmission and reception. Low latency is crucial for real-time applications like voice, video, gaming, and autonomous systems.
EVM Deviation between actual and ideal signal constellation points. Lower EVM indicates better modulation accuracy and signal quality.
II-B Sensing
Sensing refers to the process of detecting or measuring physical properties, objects, or conditions in the environment using various sensors. Sensors convert physical phenomena into measurable signals, which can be processed and analyzed. Sensing can be performed using different types of sensors, such as optical sensors, thermal sensors, acoustic sensors, and electromagnetic sensors.
Sensing has a central role in enabling data-driven decision-making, enhancing efficiency, improving safety, and advancing technological capabilities across various industries and everyday life. Application areas of sensing include, but are not limited to, the following:
  • IoT: Enabling interconnected devices to collect and share data for smart home automation, agriculture, transportation, and more.
  • Autonomous Systems: Enabling vehicles, drones, and robots to perceive and interact with their surroundings using sensors such as radar, light detection and ranging (LiDAR), cameras, and ultrasonic sensors.
  • Energy Management: Optimizing energy use and efficiency in smart grids, renewable energy systems, and smart homes.
  • Natural Disaster Monitoring: Early detection and warning systems for earthquakes, tsunamis, floods, and wildfires using sensors to mitigate risks and improve disaster response.
  • Biomedical Sensing: Monitoring and diagnosing medical conditions through sensors that detect biological signals, analyze biomarkers, and deliver targeted therapies.
As communication standards is evolving to 6G, the integration of advanced sensing technologies is expected to drive innovations in communication and computing.
II-B1 Sensing Performance Metrics
The estimation-theoretic metrics are typically used to characterize the sensing performance. In parameter estimation, the performance of the estimator is commonly evaluated using the MSE metric as $\mathrm{MSE}_{\bm{\theta}}=\mathbb{E}[\|\bm{\theta}-\hat{\bm{\theta}}\|^{2}]$, where $\bm{\theta}$ is a vector of deterministic (true) parameters and $\hat{\bm{\theta}}$ is its estimate. The estimator that minimizes the MSE is termed an MMSE estimator, and it may not be realizable for many practical scenarios. From a practical perspective, we constrain the estimator’s bias to be zero to construct an estimator that minimizes the variance. Such an estimator is termed a minimum variance unbiased estimator and its MSE is just the variance. The variance of any unbiased estimator is lower-bounded by the CRLB; $\mathrm{CRB}=\mathbf{I}^{-1}(\bm{\theta})$, where $\mathbf{I}(\bm{\theta})$ is the Fisher’s information matrix. The CRLB is widely used as a sensing performance metric. The notion of CRB can also be extended to random parameters with a known a-priori distribution. Hence, one can define a posterior CRB for Bayesian estimators as it serves as a lower bound for Bayesian MSE. Moreover, Weiss-Weinstein and Ziv-Zakai bounds are two tighter Bayesian lower bounds that can be derived when the prior distribution of the parameter vector is known.
II-C Radar
Radio detection and ranging, or radar, is a type of sensing technology that uses radio waves to detect and determine the distance, speed, and other characteristics of objects. By transmitting waves towards an object and receiving the reflected signals, radar can operate in an active sensing manner. The time delay and frequency changes (Doppler effect) between transmission and reception provide information about the object’s properties. Radar is widely used in various applications.
{forest}
[…]
Applications [air traffic control] [automotive safety systems] [weather forecasting] [military surveillance] [motion detection & perimeter security ] [fall detection and healthcare ] [security and surveillance] ]
Radars can be classified based on various criteria, including functionality and operational configuration. Common classifications include primary versus secondary, active versus passive, and monostatic, bistatic, or multistatic systems, among others.
Primary radar operates by emitting radio waves towards an object and detecting the echoes reflected back from the object. It is suitable for basic detection but provides limited additional information. Secondary radar also emits radio waves towards an object, typically an aircraft, but requires the aircraft to carry a transponder that responds to the radar signal. This ‘cooperative’ interaction provides detailed information beyond basic detection, including aircraft identification and altitude.
Bistatic, monostatic, and multistatic radars are configurations based on the relationship between the transmitter and receiver elements. Monostatic radars have both transmitter and receiver colocated, simplifying design but limiting stealth detection. Bistatic radars separate these elements, improving stealth detection and accuracy but complicating synchronization. Multistatic radars use multiple transmitters and/or receivers across different locations, enhancing range and resistance to jamming but requiring complex coordination.
A common classification of radars is given in Fig. 2. Active radar systems emit their own electromagnetic waves towards a target and detect the reflections that bounce back from objects. In contrast, passive radar systems do not emit their own waves but rely on external sources of illumination for detection. They detect reflections of signals sent from broadcast stations, communication satellites, other radars, etc. While this method allows for covert operation and reduces the risk of detection by potential adversaries, passive radars face challenges related to signal availability, processing complexity, and operational limitations compared to active radar systems.
Radaractive radar continuous-wave radar unmodulated modulated PMCW radarFMCW radarlinearnonlinearpulsed radarMTI radarDoppler radarpassive radar
Figure 2: A classification of different types of radars.
Pulsed radar works by sending short bursts of radio waves and detecting the echoes from targets. It transmits briefly and listens during the intervals between pulses, thus not listening while transmitting. The time delay between transmitted and received signals determines target distance, and the frequency shift due to the Doppler effect gives the relative speed. Pulsed radars are classified into two main types: moving target indicator (MTI) radar and Doppler radar. MTI radar is focused on filtering out stationary clutter to detect moving targets, but it does not measure target velocities as precisely. Doppler radar is designed to measure the velocity of moving targets with high accuracy. Impulse radar is a variant of pulsed radar. While both use electromagnetic pulses for detection, impulse radar operates with shorter pulses, making it ideal for high-resolution and precise distance or profile measurements.
A continuous wave (CW) radar, in contrast to pulsed radars, transmits continuously, meaning that it listens for target returns simultaneously with transmitting. CW radars can beted or unmodulated. Two prominent examples of modulated CW radars are frequency-modulated continuous wave (FMCW) and phase-modulated continuous wave (PMCW) radars. CW radar offers a longer range and higher resolution than pulsed radar but faces challenges like requiring separate transmitters and receivers, difficulty measuring target distance, and susceptibility to interference and jamming. A detailed comparison of pulsed radars and CW radar systems are given in Table III, while Table IV discusses the advantages and disadvantages of different radar types in general.
TABLE III: Comparison of pulsed radar and Continuous-Wave Radar
Feature Pulsed Radar Continuous-Wave Radar
Operation Transmits short-duration pulses of radio frequency (RF) energy. Simultaneously transmit and receive a continuous wave of RF energy.
Functionality Measures time delay for target range. Measures Doppler shift for target velocity.
Applications Air traffic control, weather monitoring, military surveillance. Speed measurement (police radar), automotive radar (adaptive cruise control).
Antenna usage The same antenna can be used for both transmission and reception (via a duplexer). Typically requires separate transmit and receive antennas.
Advantages Good range resolution, long-range detection capability. High Doppler resolution, simpler design.
Disadvantages Limited Doppler resolution, complexity in pulse compression. Limited range information without additional techniques (e.g., FMCW radar).
TABLE IV: Comparison of Advantages and Disadvantages of Different Radar Types
Type Advantages Disadvantages
Pulsed Radar Simple to implement Good for long-range detection Measures both range and speed Limited range resolution High power consumption Potential dead zones
CW Radar Measures velocity well Lower power consumption Smaller hardware Not good for simultaneous range and velocity measurement Needs a separate transmitter and receiver
FMCW Radar Measures both range and velocity Higher accuracy Low power consumption Complex signal processing Susceptible to interference
Pulse Doppler Radar High precision and resolution in speed and distance measurement Distinguishes stationary and moving targets Operates in multiple modes High power consumption High complexity Vulnerable to deception and jamming
II-C1 Radar Performance Metrics
Typical radar and sensing metrics are summarized in Table V. Detecting targets and estimating parameters are two of the primary functionalities of a radar system. In radar detection, the probability of detection, probability of false-alarm, and resolution are the widely used performance metrics. The range resolution of radar systems is defined as $\Delta R=c/2W$, where $c$ is the speed of light and $W$ is the bandwidth. Similarly, the velocity resolution is defined as $\Delta v=\lambda/(2NT)$, where $\lambda$ is the wavelength, $N$ is the number of coherently processed pulses, and $T$ is the pulse repetition interval. Finally, the half-power angular resolution is defined as $\Delta\theta=0.886\lambda/D_{A}$, where $D_{A}$ is the size of the array’s aperture. The range, velocity/Doppler, and angular accuracy can be defined by using $\Delta R$, $\Delta v$, and $\Delta\theta$.
The range (or time delay) is the only unambiguous parameter that early radar systems could estimate. The range rate (or Doppler) can be estimated by CW radars. The range and Doppler can be estimated simultaneously by pulse–Doppler radars. The angular parameters can be estimated by radar arrays. The trade-offs among range, Doppler, and angular resolutions can be characterized using the multidimensional generalized ambiguity function [38]. Under the narrowband assumption, this function can be factorized into the product of the range-Doppler ambiguity function and the azimuth-elevation angle ambiguity function. The classical ambiguity function for narrowband waveforms was first introduced to understand the resolution properties of time delay and Doppler in [39, 40].
In the seminal works [39, 40], the MSE metric between a known radar waveform $s(t)$ and its frequency-shifted and time-delayed version $s(t-\Delta\tau)e^{-j2\pi\Delta f_{d}t}$ is computed as
[…]
where $\Delta\tau$ is the time delay and $\Delta f_{d}$ is the frequency shift. The only term in (1) that depends on $\Delta\tau$ and $\Delta f_{v}$ has been defined as the classical ambiguity function
[…]
where $s^{*}(t)$ denotes the complex conjugate of $s(t)$.
This two-dimensional ambiguity function can also be interpreted as the matched filter output in the receiver when a delay mismatch $\Delta\tau$ and a Doppler mismatch $\Delta f_{d}$ occur. The value of $\chi(\Delta\tau,\Delta f_{d})$ evaluated at $\Delta\tau=0$ and $\Delta f_{d}=0$ corresponds to the matched filter output without any delay-Doppler mismatch. Thus, the larger the ambiguity function at the origin, the better the time-delay (range) and Doppler resolution. Moreover, the ambiguity function must be much smaller for $\Delta\tau\neq 0$ and $\Delta f_{d}\neq 0$. This is because the correlation between $s(t-\Delta\tau)e^{-j2\pi\Delta f_{d}t}$ and $s(t-\Delta\tau^{\prime})e^{-j2\pi\Delta f_{d}^{\prime}t}$ corresponds to $\chi(\Delta\tau-\Delta\tau^{\prime},\Delta f_{d}-\Delta f_{d}^{\prime})$. There exists a fundamental ambiguity or duality between target resolution in time and frequency. The targets cannot be resolved perfectly in time and frequency simultaneously for a given time–bandwidth product, and this is the reason for its name; the ambiguity function.
TABLE V: Radar Performance Metrics for Radar Systems
Metric Definition Description & Additional Information
Range resolution Ability to distinguish closely spaced targets. Higher resolution allows for better target discrimination.
Doppler resolution Ability to distinguish targets by velocity. Important for detecting moving targets.
Ambiguity function The resolution and ambiguity in range and Doppler. Helps analyze the trade-off between range and Doppler resolution.
Probability of detection ($P_{D}$) Likelihood of correctly detecting a target. Higher $P_{D}$ is desirable for reliable detection.
Probability of false alarm ($P_{\rm FA}$) Likelihood of false target detection. Lower $P_{\rm FA}$ is desirable to reduce false alarms.
ROC $P_{\rm D}$ versus $P_{\rm FA}$ for a given decision threshold. ROC curve evaluates a system’s ability to detect the presence of a signal, target, or object versus making a false alarm.
Clutter rejection Ability to distinguish targets from clutter. Effective clutter rejection improves target detection.
Azimuth/elevation resolution Ability to distinguish targets by angular position. Higher angular resolution allows for better spatial discrimination.
CRLB Theoretical lower bound on the variance of unbiased parameter estimators. Provides a benchmark for the best achievable estimation accuracy (e.g., range, velocity).
SNR Ratio of target signal power to noise power. Higher SNR improves detection performance.
MSE Average MSE between estimated and true parameter values. Measures the accuracy of parameter estimation (e.g., range, velocity). Lower MSE indicates better estimation performance.
II-D Integrated Sensing and Communication
In its broad sense, ISAC refers to the integration of various sensing technologies, including radar, with communication systems. The goal is to leverage shared resources, infrastructure, or information exchange between sensing (like environmental monitoring, surveillance) and communication functions (like data transmission, networking). It often focuses on optimizing sensor data utilization within communication networks or enhancing sensing capabilities through communication systems.
However, ISAC typically denotes the specific integration of radar systems with communication systems. This involves sharing hardware components, signal processing resources, or network infrastructure between radar functions (such as target detection, tracking) and communication functions (data transmission, networking). The emphasis is on synergizing radar functionalities with communication technologies to improve efficiency and effectiveness in both domains. In this paper, we focus on ISAC within this specific domain.
Integrating radar and communication systems poses several significant challenges. Achieving the diverse operational requirements and technical specifications of radar and communication systems is one main challenge. Radar systems typically prioritize high-resolution target detection and tracking capabilities. On the other hand, communication systems focus on efficient data transmission, network scalability, and reliability. Integrating these systems requires addressing conflicting requirements, optimizing resource allocation, and ensuring seamless interoperability without compromising performance in either domain.
II-D1 ISAC Performance Metrics
The performance of ISAC designs can be evaluated by characterizing fundamental trade-offs between sensing and communication metrics [5, 41, 42]. Generally, an ISAC metric combines communication and radar matrices from Tables II and V, along with newly defined notions.
The concept of radar estimation information rate, introduced in [41, 43], parallels the data information rate in communication systems. It treats radar illumination on a target as a form of non-cooperative communication, where parameters such as delay, Doppler, and spatial directions are transmitted. Consequently, the radar channel can be viewed as a non-cooperative communication channel. This information rate is defined as the mutual information between radar and target, derived from the entropy of random estimation parameters and the entropy of estimation uncertainty [41]. Using the data processing inequality, the mutual information between the observation $Y$ and the true parameter $\theta$ is lower-bounded as [34] $I(Y;\theta)\geq\frac{1}{2}\log{\frac{P}{D}},$ where $\theta$ is Gaussian-distributed with variance $P$, and $D$ is the MSE distortion of estimating $\theta$. This bound links the equivalent estimation information rate to the MSE distortion.
Alternatively, an equivalent communication information rate was introduced in [43], leveraging rate-distortion theory to relate MSE distortion $D$ to the information rate $R$ as $D=2^{-R}$ assuming a Gaussian channel. The trade-off between sensing distortion and communication capacity can also be characterized via the capacity-distortion function [34]. Furthermore, fundamental trade-offs among achievable sum rate, detection probability, and CRLBs can be quantified to assess ISAC performance [42].
III Multi-Objective Optimization: Traditional versus Learning-Based Approaches
ISAC problems are inherently multi-objective optimization problems. Such a problem involves simultaneously optimizing two or more conflicting objectives which can be formulated as
$\underset{}{\text{Minimize}}$ $\quad\bm{f}(x)=[f_{1}(x),f_{2}(x),\ldots,f_{i}(x)]$ (3) Subject to $\quad{g}_{j}(x)=0,\quad j=1,2,\ldots,m$ $\quad{h}_{k}(x)\leq 0,\quad k=1,2,\ldots,n\quad$
in which $f_{i}(x)$ are the objective functions, $g_{j}(x)$ and $h_{k}(x)$ are the equality and inequality constraints, respectively, and $x$ is the decision variable vector.
There are several approaches to solving multi-objective optimization problems. In this section, we will discuss these approaches, classifying them to traditional methods and learning-based techniques.
III-A Traditional Methods
Using the scalarization method [44], we can transform a multi-objective optimization problem into a single-objective problem. This can be done in various ways such as weighted sum method and $\epsilon$-constraint method, as explained below.
III-A1 Weighted Sum Method
It is also known as linear scalarization. The design is equivalent to optimization of a joint cost function that balances the trade-off between communication and radar sensing objectives (optimizing the corresponding performance metrics). It can be described as the following optimization problem.
$\underset{}{\text{Minimize}}$ $\quad\eta\,f_{\rm sens}(\cdot)+\bar{\eta}\,f_{\rm comm}(\cdot)$ (4) Subject to $\quad{g}(\cdot)=0,\quad$ $\quad{h}(\cdot)\leq 0,\quad$
where $\eta$ ($0\leq\eta\leq 1$) and $\bar{\eta}=1-\eta$ are the normalizing and weighting factors assigning the priorities for sensing and communication tasks, respectively, and $f_{\rm sens}({\rm sensing\;metric})$ and $f_{\rm comm}({\rm communication\;metric})$ are some functions of sensing and communication metrics, respectively. The weights can be adjusted adaptively with respect to the requirements imposed by different scenarios. Various sensing and communication metrics can then be used.
Only some simplified forms of the problem above under some simplifying assumptions, which may not hold in practice, can be solved using optimization-based approach.
III-A2 $\epsilon$-Constraint Method
In this method, one objective is optimized, while the other objectives are treated as constraints with acceptable bounds ($\epsilon$). By adjusting the $\epsilon$-values, different Pareto-optimal solutions can be generated. This method is particularly useful when there is a priority among objectives as it prioritize one objective while constraining the others.
Mathematically, the optimization can be expressed as
$\text{Minimize}\quad f_{i}(\cdot)$ Subject $\text{Subject to}\quad f_{j}(\cdot)$ $\leq\epsilon_{j},\quad j\in\{{\rm sens},{\rm comm}\}\backslash\{i\}$ ${g}(\cdot)$ $=0,\quad$ $\quad{h}(\cdot)$ $\leq 0,\quad$ (5)
where
  • $f_{i}(\cdot)$ is the primary objective function to minimize;
  • $f_{j}(\cdot)$ are the other objective functions, constrained by the thresholds $\epsilon_{j}$; and,
  • $\epsilon_{j}$ are user-defined constraint bounds.
The $\epsilon$-constraint method can generate solutions on non-convex parts of the Pareto front, which the weighted sum method may miss. Also, unlike the weighted sum method, it does not require selecting weights for different objectives, thus avoiding the potential bias introduced by weight choices. This makes the $\epsilon$-constraint method more flexible and suitable for problems where it is challenging to assign relative importance to objectives.
III-A3 Pareto Optimization
This approach, instead of finding a single optimal solution, aims to find the Pareto front—a set of solutions where no objective can be improved without degrading another [45, 46]. This can be achieved using techniques like evolutionary algorithms and Pareto-based methods (see, for example [47]).
III-B Learning-Based Techniques
ML-based approaches can be divided into supervised, unsupervised, and reinforcement learning, in general. Each approach can be implemented in various ways, such as multi-objective versus single-objective learning, multi-task versus single-task learning, and end-to-end learning, or a combination of them. In this section, we first elaborate on multi-objective and multi-task learning, highlighting their differences. We then cover end-to-end learning and hybrid methods.
[…]
Figure 3: Comparison of single-task and multi-task learning for ISAC systems. Left: single-task learning trains separate models for each task independently. Right: Multi-task learning leverages shared layers to learn common features before branching into task-specific layers.
III-B1 Multi-Task Learning
Multi-task learning (MTL) is an ML paradigm where a model learns to perform multiple tasks simultaneously. Each task has its own set of outputs, and the model learns to predict those outputs jointly. The main idea is that learning shared representations across tasks can improve generalization and performance on each individual task. In contrast, single-task learning focuses on training a model for one specific task at a time, without considering interactions or dependencies with other tasks. Single-task learning is suitable when the goal is to solve a single, well-defined problem without considering the interactions or dependencies with other tasks. It is commonly used when the tasks are distinct and unrelated, or when there is insufficient data or computational resources to train a multi-task model. In multi-task learning, the goal is to train a model to perform multiple tasks simultaneously. Figure 3 illustrates the difference between these two types of learning.
The model combines these task-specific losses into a single joint loss function, usually by applying weighted sums
$\mathcal{L}_{\text{MTL}}=\sum_{i=1}^{T}\lambda_{i}\mathcal{L}_{i}$ (6)
where $\mathcal{L}_{i}$ is the loss for task $i$, $\lambda_{i},\;i=1,\cdots,T$ are weights balancing the importance of the tasks, and $T$ is the number of tasks. The overall objective is to minimize the joint loss across all tasks.
Example: Consider a self-driving car neural network model that has two tasks: detecting pedestrians (object detection) and predicting the road’s curvature (regression). These are two distinct tasks, and the model combines the losses from both tasks to optimize neural network’s performance, ensuring that it effectively detects pedestrians while accurately predicting road curvature
$\mathcal{L}_{\text{MTL}}=\lambda_{1}\cdot\mathcal{L}_{\text{classification}}+\lambda_{2}\cdot\mathcal{L}_{\text{regression}}.$ (7)
Goal: The goal is to optimize performance on multiple tasks using shared information or features.
The core idea of multi-task learning is that learning related tasks together improve performance by allowing the model to share information and transfer knowledge across tasks. This is especially effective when tasks have common features or underlying similarities.
Multi-task learning is implemented using either hard or soft parameter sharing approaches [48], explained below.
  • In hard parameter sharing, a subset of parameters is shared across all tasks, while the remaining parameters are task-specific. This approach reduces the risk of overfitting by utilizing shared information, making it particularly effective when tasks are closely related.
  • In soft parameter sharing, each task maintains its own task-specific parameters, but these are jointly regularized to encourage knowledge transfer between tasks. This is commonly achieved through techniques like Bayesian priors or a shared dictionary, allowing the model to learn task relationships without enforcing direct parameter sharing.
III-B2 Multi-Objective Learning
Similar to multi-objective optimization, multi-objective learning (MOL) aims to optimize multiple, possibly conflicting, objective functions simultaneously. These objectives represent different criteria that need to be balanced within the same task. The goal is to find a balance between these objectives, often resulting in a set of solutions known as the Pareto front.
In multi-objective learning, the model optimizes multiple objectives within a single task. Each objective has its own loss function, but instead of combining them into a single joint loss, the goal is to balance or trade off between the competing objectives. A common approach is to use weighted sum scalarization to combine the objectives into a single scalar loss function
$\mathcal{L}_{\text{MOL}}=\sum_{i=1}^{O}\alpha_{i}\mathcal{L}_{i}$ (8)
where $\mathcal{L}_{i}$ is the loss for objective $i$, $\;i=1,\cdots,O$, $\alpha_{i}$ are weights, and $O$ is the number of objectives. Alternatively, other methods like the $\epsilon$-constraint method or Pareto optimization can be used to solve the multi-objective optimization problem without explicitly combining the losses into one scalar function.
Example: Consider a single-task car model trained solely to detect pedestrians. This could still have multiple objectives, e.g., one objective might be to maximize accuracy, while another is to minimize inference time. These are two conflicting objectives for the same task, requiring the model to find an optimal trade-off between them, e.g., by
$\mathcal{L}_{\text{MOL}}=\alpha_{1}\cdot\mathcal{L}_{\text{accuracy}}+\alpha_{2}\cdot\mathcal{L}_{\text{latency}}.$ (9)
Goal: The goal is to optimize multiple ‘conflicting’ objectives simultaneously. The model does not simply combine the losses but tries to balance them, often leading to Pareto optimal solutions where improving one objective worsens another.
Remark 1.
The key difference between a task and an objective is that a task refers to what the model is designed to do, such as detecting objects, predicting the road’s curvature, or removing interference. An objective, on the other hand, defines how the model’s performance is measured and optimized, such as minimizing error, maximizing accuracy, or minimizing latency. In multi-task learning, the model is trained to solve multiple tasks simultaneously, whereas in multi-objective learning, the focus is on optimizing multiple objectives for a single task.
In summary, multi-task learning is about performing multiple tasks simultaneously and using shared representations to improve performance. Multi-objective learning is about balancing multiple conflicting objectives during the optimization process.
Remark 2.
It is natural to treat sensing and communication as two distinct tasks, making ISAC problems inherently multi-task learning problems. However, multi-task learning problems can also be approached through multi-objective optimization [48]. In addition, in traditional multi-objective optimization, sensing and communication are considered as separate objectives, blurring the line between multi-task learning and multi-objective optimization in the context of ISAC. As such, several works, such as [49, 50, 3], have treated sensing and communication as distinct objectives.
IV Waveform and Beamforming Design
In this section, we cover the basic definitions, required criteria, and common types and examples for waveforms and beamforming in radar, communication, and ISAC.
IV-A Waveform Basics
A waveform refers to the shape or structure of the electromagnetic signal that is transmitted over a medium (channel) for communication, radar, ISAC, or other purposes. It describes how the signal varies over time, typically in terms of its amplitude, frequency, or phase. While the core definition of a waveform is consistent across radar and communication systems, the objectives and purposes behind transmitting the waveform differ largely between the two domains, as do their performance criteria.
IV-A1 Waveform in Radar
In radar systems, the waveform is designed primarily for sensing and detecting objects. A well-designed waveform can greatly enhance system performance by improving the signal-to-interference ratio, optimizing spectrum usage, aligning with desired beampatterns, and improving the accuracy of target parameter estimation, among other factors [51, 52, 53].
Waveform design in radar involves selecting specific parameters such as pulse duration, pulse repetition frequency, carrier frequency, modulation scheme, and waveform shape (e.g., pulsed, continuous wave, frequency-modulated) to achieve desired performance objectives. These objectives may include maximizing detection range, improving target discrimination, and minimizing the radar’s detectability. The choice of waveform can significantly impact radar system performance in terms of range, accuracy, and sensitivity to various types of targets and interference. For example, short-duration pulses may be used for high-resolution imaging radar, while longer pulses or continuous waveforms might be employed for detecting moving targets or tracking objects over long ranges. Common radar waveforms are [54]
  • Pulsed Waveforms: The radar sends out short bursts of high-power energy.
  • Chirp Waveforms: A linear frequency-modulated waveform, where the frequency of the signal changes linearly over time. Chirps are widely used because they provide high resolution in range and are robust to noise.
  • Continuous Waves: They involve continuous transmission without interruptions and can be unmodulated or modulated. The former is used for velocity measurement via the Doppler effect, while modulated CW radars—such as FMCW and PMCW—enable range estimation.
  • OFDM Waveforms: Originally developed for communication systems, OFDM waveforms have been adapted for radar because of their flexibility in spectrum shaping and potential for joint radar-communication applications.
IV-A2 Waveform in Communication
In communication systems, a waveform serves as the carrier of information from a transmitter to a receiver, encoding data into a format that can be efficiently transmitted over a communication channel and decoded at the receiver. Waveform design in communication involves selecting parameters such as modulation scheme, symbol rate, carrier frequency, bandwidth, and pulse shape to optimize efficiency and reliability of the transmission for the intended application and environmental conditions. OFDM has been the primary waveform used in cellular networks starting from 4G.
IV-A3 Waveform in ISAC
In ISAC, a single unified waveform is used to accomplish both sensing and communication tasks. Communication waveforms can be directly employed for sensing. The multicarrier waveforms based on OFDM have shown good performance in both sensing and communication without requiring substantial alterations to existing communication system designs [55], [56]. Other waveforms such as OTFS [57] and orthogonal delay-Doppler division multiplexing [58] are shown to better suit to time-varying channels, especially in high-mobility scenarios. For example, OTFS shows greater robustness to Doppler shifts compared to OFDM in communication scenarios [59], while both waveforms exhibit comparable performance for sensing applications [60]. The difficulty with such communication-centric waveforms is that their peak-to-average power ratio (PAPR) is greater than one, indicating that the envelope of the transmitted time-domain signal does not have a constant amplitude. For example, the expected PAPR of OFDM increases with the number of subcarriers, while the expected PAPR pf OTFS increases with the number of subsymbols transmitted over slow-time in one radar frame. As a result, a non-uniform envelope requires using highly linear power amplifiers at the transmitter to avoid signal distortion that affects both radar and communication performance. It conflicts radar requirements, since high power is always beneficial for sensing. Indeed, the attenuation of a signal power for both way propagation in sensing channel is bigger than that for one way propagation in communications channel. OTFS suffers from this drawback in ISAC, just like OFDM. To deal with such difficulties, AI-based solutions have shown their importance within this approach based on the use of communication-centric waveforms [61, 62].
A robust waveform design method for multiple-input and multiple-output (MIMO) dual-functional radar-communication (DFRC) system in the presence of transmit hardware impairments has been proposed in [63]. In [64], the authors devised a power-domain non-orthogonal ISAC waveform design with OFDM modulation, eliminating the need for complex operations while providing flexible control over the trade-offs between sensing and communication performances. In [65], to address the problem of short sensing range achieved by radar sensing in cyclic prefix OFDM based ISAC system, a novel reference signal design and corresponding signal processing scheme were developed for OFDM long-range radar sensing. Relevant to waveforms design other ISAC functionalities are sidelobe information embedding in radar systems [66], coexistence through array partitioning and waveform optimization [67], or index modulation through exploitation of the different degrees of freedom in radar systems [68].
A generic waveform structure that allocates a fraction of symbols to communication (used exclusively for data transmission) and the remaining fraction to preamble symbols (dedicated entirely to sensing) within a coherent processing interval of $T$ seconds spanning $Q$ frames can mitigate the major limitations associated with communication-centric waveforms. Such generic waveform structure has been introduced for ISAC, especially for systems with high requirements to accuracy of velocity estimation, in [5, 69]. The IEEE 802.11ad standard can support such a multi-frame approach, where, as illustrated in Fig. 4, each frame consists of a fixed preamble duration $P$ and a variable data payload, resulting in a varying overall frame length. The preambles are used as radar waveforms (e.g., pulsed waveforms), which are distinct from the communication waveforms (e.g., OFDM-based waveforms) employed for data transmission.
Figure 4: A coherent processing interval of $T$ seconds duration with $M$ ISAC frames. Each frame contains a fixed-length preamble of $P$ symbols and a varying length data segment. The length of each frame is an integer multiple, $q_{m}$, of the Nyquist sampling interval in the Doppler domain, $T_{\rm D}$.
Using the frame structure in Fig. 4, the general ISAC performance optimization problem that aims to achieve the best possibly sensing performance by using the least possible resources (number of preambles in Fig. 4) can be formulated as
$\underset{Q,\{m_{q}\}_{q=1}^{Q-1}}{\text{Minimize}}$ $\eta\,f_{\rm sens}(Q,\{m_{q}\}_{q},\cdot)+\bar{\eta}\,f_{\rm comm}(Q,\{m_{q}\}_{q},\cdot)$ (10)
Subject to $\{T,K\}=\mathrm{constants,}$
$0<m_{1}<\cdots<m_{Q-1}<T/T_{\rm D},$
where $\eta$, $\bar{\eta}$, $f_{\rm sens}(Q,\{m_{q}\}_{q=1}^{Q-1},\cdot)$ and $f_{\rm comm}(Q,\{m_{q}\}_{q=1}^{Q-1},\cdot)$ have been introduced after (4), and $K$ is the number of targets to be sensed. Alternatively, problem (4) can be modified as a minimization of one of the objectives with the second one written as an $\epsilon$-constraint (see (III-A2)) that would guarantee an acceptable performance for one of the tasks. Such problem simplifies whenever a specific sparse pulse configuration/preamble positions, such as coprime pulses, is assumed. However, in general, the problem of optimal preamble location finding is hards (similar to a sphere packing problem), and there is no efficient optimization-based solution to it. It may however be addressed using AI methods. Thus, despite not being fully explored yet, AI can serve as an enabling approach for ISAC waveform design, when exact mathematical or optimization-based approaches fail.
IV-B Beamforming Design
Beamforming, or precoding, the control of signal phase and amplitude, is a key signal processing technique used in both wireless communications and radar. While both applications share the same fundamental principles, their goals and implementations differ.
  • In wireless communications, beamforming directs signals toward intended users, minimizes interference, enhances spectral efficiency, extends coverage, and improves overall system performance. Its focus is on optimizing data transmission.
  • In radar, beamforming is primarily used for target detection, localization, and resolution. It emphasizes directional sensitivity and interference rejection to improve detection accuracy and target tracking.
A review of 25 years of progress in beamforming, tracing its evolution from convex and nonconvex optimization techniques to modern machine learning approaches, with applications in radar, communications, and other domains can be found in [70]. In the following, we describe different beamforming methods used in communication, radar, and ISAC.
IV-B1 Beamforming in Communication
In wireless communications, beamforming is a critical technology enabling higher data rates, improved connectivity, and enhanced spectral efficiency. Beamforming can be implemented in digital, analog, or hybrid forms, each with its own advantages and trade-offs.
TABLE VI: Comparison of Digital Beamforming (DBF) Techniques in Communication.
DBF Tech. Description   Advantages & Disadvantages Beamforming Formula Ref.
[…]
ZF Eliminates multi-user interference by inverting the channel matrix. + Simple implementation, effective at high SNR. Amplifies noise at low SNR; performance degrades in noisy environments. $\mathbf{W}_{\rm ZF}=\mathbf{H}^{H}(\mathbf{H}\mathbf{H}^{H})^{-1}$ [3]
MMSE Balances interference suppression and noise amplification by utilizing a regularization factor $\alpha$ in the channel inversion. + Performs well at both low and high SNR, more robust than ZF. Higher computational complexity than ZF. $\mathbf{W}_{\rm MMSE}=\mathbf{H}^{H}(\mathbf{H}\mathbf{H}^{H}+\alpha\mathbf{I})^{-1}$ [4]
MRT/MRC Maximizes SNR by transmitting/receiving in the direction of the strongest channel gain. + Maximizes received power, simple to implement. Does not cancel interference; not suitable for multi-user scenarios. $\mathbf{w}_{\rm MRT}=\frac{\mathbf{h}\dagger}{\|\mathbf{h}\|}$ [74]
[…]
ML-based Uses deep learning to optimize beamforming in dynamic and complex environments. + Adapts to non-stationary channels; reduces computational overhead. Needs large datasets for training. AI-based optimization, no explicit formula [77, 78, 70]
Digital Beamforming
Digital beamforming (DBF) operates entirely in the baseband domain, leveraging digital signal processing techniques to manipulate signals amplitude/phase before transmission. Each antenna element is connected to an independent RF chain, allowing precise beam steering and adaptive beamforming capabilities. This flexibility enables multi-user communication and dynamic interference mitigation, making digital beamforming a key block of advanced MIMO systems. However, its implementation requires high-speed analog-to-digital and digital-to-analog converters, which lead to increased power consumption and hardware complexity. Despite these challenges, DBF continues to be the most adaptable and widely utilized approach for multi-user communication where fine-grained control over multiple simultaneous beams is needed.
DBF in communication uses various techniques depending on system goals, complexity, and constraints. Singular value decomposition (SVD) and eigenvalue decomposition (EVD) provide optimal performance in single-user MIMO, while zero-forcing (ZF) and MMSE beamforming are widely used in multi-user MIMO to handle interference. Maximum ratio transmission (MRT) and maximum ratio combining (MRC) are simple techniques, but they lack interference mitigation, whereas Capon beamforming, also known as minimum variance distortionless response (MVDR) beamforming, is effective for adaptive spatial filtering. With emerging ML–based sensing and beamforming, future beamforming techniques are expected to become increasingly environment-aware, adaptive, and intelligent, paving the way for ISAC and 6G systems. Common DBF techniques are summarized in Table VI.
Analog Beamforming
Analog beamforming, in contrast, controls the phase and amplitude of signals directly in the RF domain using phase shifters before converting them to baseband. This approach forms a single beam per RF chain, making it highly power-efficient and well-suited for high-frequency mmWave communications. Its lower hardware complexity makes it an attractive choice for cost-sensitive deployments, especially in systems where power consumption is the main concern. However, analog beamforming lacks the ability to generate multiple independent beams simultaneously, limiting its flexibility in supporting multi-user MIMO scenarios. The reduced control over beamforming patterns also poses challenges in dynamically adapting to varying channel conditions, making it less effective in highly dynamic environments.
Hybrid Beamforming
Hybrid beamforming compromises between digital and analog techniques, balancing flexibility and efficiency by reducing the number of RF chains [70] while maintaining some level of digital control. This approach is particularly advantageous in massive MIMO and mmWave systems, where the cost and power consumption of fully digital beamforming become prohibitive. By integrating digital processing at the subarray level and analog processing at the RF front-end, hybrid beamforming allows for multiple beam generation while optimizing energy efficiency. However, its implementation requires sophisticated optimization algorithms to effectively manage the trade-offs between beamforming precision and hardware constraints. The performance of hybrid beamforming depends heavily on antenna architecture and system design, but it represents a promising solution for next-generation wireless networks, including 5G and beyond.
IV-B2 Beamforming in Radar
Like in communication systems, beamforming in radar is used to direct and focus transmitted energy in specific directions while enhancing target detection and reducing interference. It improves performance by shaping the radiation pattern of an antenna array [79, 9, 80, 81, 82]. However, radar beamforming differs from communication beamforming in objectives, processing techniques, and system constraints. In radar, the primary goal is target detection, tracking, and imaging, while in communication, beamforming aims to enhance data transmission fidelity and spectral efficiency.
  • Conventional Beamforming: A fixed-direction technique that uses a static antenna pattern, achieved via mechanical steering (e.g., rotating parabolic dishes) or phased arrays, to direct the beam across a region.
  • Adaptive Beamforming (Digital Beamforming): Dynamically adjusts antenna weights based on received signal characteristics, optimizing the beampattern in real time for clutter suppression, interference rejection, and multi-beam processing. It is often implemented in phased array radars, provides improved resolution, tracking accuracy, and flexibility compared to fixed beamforming.
  • Analog Beamforming: Similar to communication, analog beamforming in radar relies on RF phase shifters to control signal direction. Unlike digital beamforming, which processes signals in the digital domain, analog beamforming operates directly on the analog signals.
  • Hybrid Beamforming: Combines analog and digital techniques for cost-performance balance.
  • MIMO Beamforming: Employs multiple transmit/receive antennas for improved angular resolution and target separation. It excels in low SNR conditions and is particularly useful for multi-target tracking.
  • SAR and ISAR Beamforming: Synthetic aperture radar (SAR) and inverse SAR (ISAR) beamforming synthesize a large aperture using motion to achieve high-resolution imaging. SAR is commonly used in earth observation and remote sensing, while ISAR is applied to moving target imaging, such as ships and aircraft.
IV-B3 Beamforming in ISAC
An ISAC beamformer must optimize sensing and communication accounting for trade-offs. In [83], a joint active beamforming design for base station (BS) and passive beamforming design for intelligent reflecting surfaces was proposed to enhance both communication and localization performance in the downlink scenario. Similarly, in [84], the authors studied ISAC beamforming design by maximizing the MI between the target response matrix of a point radar target and the echo signals, while ensuring the data rate requirements of the communication users. Three types of echo interference were considered: no interference, point interference, and extended interference.
Interference management has also been explored in other studies. In [85], the authors investigated the precoder optimization problem to maximize radar sensing SINR, considering both signal-independent interference and signal-dependent clutter. The proposed approach ensured a minimum SINR at multiple communication users while maintaining per-antenna power constraints. Additionally, in [86], a communication-centric subcarrier allocation strategy was developed for wideband ISAC systems, where a subset of subcarriers was tailored for radar sensing while the entire bandwidth was used for communications. An efficient hybrid beamforming design was proposed on the basis of this allocation scheme.
To improve spectral efficiency, researchers have explored full-duplex ISAC systems. For example, a full duplex ISAC system where the base station performs target detection while communicating with both downlink and uplink users using the same time and frequency resources was studied in [87]. An energy-efficient robust beamforming design for a full-duplex ISAC system under norm-bounded CSI errors was introduced in [88]. In [89], a new communication receiver design for a downlink multi-user communication and multi-target ISAC scenario was proposed which is capable of canceling interference caused by radar signals and maximizing the minimum weighted beampattern gain at the sensing angles of interest.
Robustness under imperfect CSI was the focus of [90], where the authors studied the robust DFRC beamforming design taking into account outage SINR probability under imperfect CSI. In [91], a multi-beam design scheme was proposed where one beam is used to communicate with a single-antenna user in the downlink while one or multiple dedicated sensing beams are used to sense the unknown and random angle parameter of a target. Similarly, [92] introduced a beam pattern modulation-embedded hybrid transceiver design for mmWave ISAC systems, where the ISAC transmitter provides multi beams for single-user communication and scanning beams for sensing.
In [93], transmit beamforming was studied for a multi-target, multi-user MIMO-ISAC system, where a general form and tight upper bound of sensing MI were derived, along with a multi-objective optimization framework based on user SINR and sensing MI. In [94], to mitigate interference between target echo signals while capturing near-field channel characteristics in a full-duplex ISAC system, the authors designed optimal transmit beamforming and receive beamforming, and developed an iterative beamforming algorithm to solve the resulting optimization problem. In [95], the authors exploited both the training and transmission signals for sensing, and proposed two MSE-based schemes for training and transmission design.
IV-C Joint Beamforming and Waveform Design
Joint beamforming and waveform design can be approached with various objectives in mind. Maximizing the SNR is crucial for radar applications, while minimizing different types of interference (maximizing SINR) is essential for communications quality and transmission rate. In [96], the authors designed transmit precoders and receive combiners to minimize inter-user, intra-user, and radar-communications interference for ISAC systems. In [4], considering point and extended target scenarios, the authors proposed minimizing the CRB of radar sensing while ensuring a predefined level of SINR for each communication user.
Constellation design is another related topic. The waveform and constellation are closely related as constellation serves as the “alphabet” of a modulation scheme, defining the set of possible symbols, while the waveform is the “spoken sentence” that delivers those symbols over the air. In OFDM systems, high-rate communication prefers constellations with varying amplitudes—such as QAM—resulting in non-constant amplitude waveforms with higher PAPR but improved spectral efficiency and error performance. In contrast, constant-modulus constellations like PSK yield waveforms with lower PAPR and better auto-correlation properties, making them more suitable for sensing tasks such as radar. Therefore, ISAC systems, waveform and constellation design must strike a balance between the randomness required for communication and the determinism favored by sensing [97]. Moreover, constellation structure plays a key role in enabling effective interference cancellation.
V ML-based Waveform and Beamforming Design for ISAC
We describe different ways to optimize (3), or its variants in (4) and (III-A2), using learning-based techniques in what follows.
V-A Single-Task Learning-Based Techniques
Many of the existing works on ISAC fall in this category.
Single-task learning aims to train a network to optimize $\eta\,f_{\rm sens}(\cdot)+\bar{\eta}\,f_{\rm comm}(\cdot)$ as one task where sensing and communication objectives are optimized concurrently. In this approach, $\eta$ could be an input parameter in addition to other inputs such as channel coefficients. The weight could also be fixed if there is a preference for one of the objectives. This itself could be performed using supervised, unsupervised, and reinforcement learning. Many of the existing works on ISAC fall in this category.
V-A1 Supervised Learning
In [98], a graph neural network (GNN)-based binary classifier was designed to determine irrelevant nodes, thereby reducing computational complexity of branch and bound algorithm tailored to the general multi-user transmit beamforming design problem for ISAC systems. In [99], a framework for sensor-aided beam training in a multi-user mmWave MIMO system is proposed. It designs an ML–based multimodal fusion network trained on data from camera, LiDAR, and position sensors, along with the beamspace representations of the channel. In [100], a model-based ML approach was proposed for ISAC receiver design using a simplified sliding transformer. Li et al. [101] proposed a GNN-based approach using heterogeneous graph representation to jointly optimize service mode selection and target association in THz vehicular networks, aiming to maximize communication data rates while meeting sensing requirements. In [102], a model-driven regression network combining a multi-layer neural network and a single-layer perceptron is proposed to enhance near-field emitter localization accuracy.
V-A2 Unsupervised Learning
Several studies have leveraged deep learning to enhance predictive beamforming. To enhance hybrid beamforming, [103] proposed a modified projected gradient ascent method and an interpretable deep unfolded projected gradient ascent framework. In [104] a convolutional long short-term memory (LSTM) network is proposed that uses historical channels for predictive beamforming in ISAC vehicular networks, while [105] uses a neural network for statistical joint optimization of beamforming and localization based on environmental data from near-field channel estimation.
In [106], a joint design of sensing waveform and communication receive beamforming was proposed to mitigate mutual interference in a non-orthogonal uplink ISAC system, with an unsupervised deep neural network (DNN) structure. Using deep learning to learn key parameters of the weighted MMSE, [107] proposed manifold optimization for hybrid precoding in mmWave multiuser MIMO systems. In [108], the authors addressed Gaussian channel uncertainties by developing a modified optimal beamforming structure and a specialized GNN for bipartite message-passing inference.
In [109], a deep unrolling approach was introduced to design downlink hybrid MIMO precoders with speed, robustness, and interpretability. Similarly, [110] proposed a beamforming learning architecture by unfolding a parallel gradient projection algorithm. Moreover, generative adversarial networks (GANs) have been applied to ISAC tasks; for example, [61] utilized GANs to adaptively generate jamming waveforms based on specific jamming effect requirements. In [111] an ISAC waveform design based on unsupervised learning is proposed to address the high complexity of classical optimization, using a custom loss function that balances sensing and communication trade-offs. It demonstrats competitive performance with reduced complexity which is especially suitable for large antenna arrays.
V-A3 Reinforcement Learning
In [62], the authors introduced a model-based reinforcement learning framework for waveform optimization in multi-carrier ISAC systems operating under dynamic environments, offering enhanced explainability and sample efficiency. In [112], a constrained deep reinforcement learning approach was proposed for real-time precoding in ISAC systems, enabling simultaneous target tracking and multi-user communication. To improve training efficiency, the method integrates a primal-dual deep deterministic policy gradient algorithm with the Wolpertinger architecture. Most reinforcement learning-based ISAC schemes focus on simple point-to-point channels, and their extension to multi-user and multi-cell [113] scenarios are unexplored.
V-B Multi-Task Learning
While ISAC problems can be addressed using either single-task or multi-task learning, the latter is generally more suitable. As an alternative to single-task learning, we may train a single neural network for multiple task individually. Specifically, we may train the network to generate a waveform for $f_{\rm sens}(\cdot)$ and then retrain it for $f_{\rm comm}(\cdot)$. In this approach, a binary vector can be used as an input to distinguish the tasks, like in [114]. Even the radar component itself can involve multiple tasks. Traditional radar processing is modular, consisting of separate stages such as signal detection, target tracking, and classification. More specifically, radar processing typically involves sequential stages:
  • Signal Processing: Tasks such as filtering, detection, and parameter estimation (e.g., range and velocity).
  • Tracking: Using algorithms like Kalman filtering to track detected targets.
  • Classification: Identifying target types based on additional signal features.
Each stage is independently designed and optimized, which can lead to suboptimal overall performance. However, deep learning—particularly multi-task learning—integrates all these stages into a single model that is optimized jointly, from raw data acquisition to final decision-making, for the desired outcomes. Similarly, the communication component itself may involve multiple tasks.
V-C Model-Based Learning
In both sensing and communications, we often have access to system or process models. However, even when these models are adopted in corresponding optimization problems, acceptable complexity of optimization-based solutions is not guaranteed. In addition, optimization-based algorithms have to be typically manually designed and allow for few or no tuning parameters. Then the model-based learning or learning-to-optimize approach is preferable. Using the model, the learned optimizer is trained over a set of similar optimizees and as a result is able to solve unseen optimizees from the same distribution [115]. The following two approaches are typically used to design such learned optimizers.
V-C1 Plug-and-Play
An optimization algorithm may be too complex. In such a case, the plug-and-play approach can be applied to replace the most computation-intense components with a neural network [116], which is then trained on a set of similar optimizees, as previously described. This approach is especially useful when an optimization algorithm can be decomposed into update rules, allowing computationally intensive steps to be replaced with learnable neural networks.
V-C2 Algorithm Unrolling
Algorithm unrolling is an alternative approach for model-based learning. It consists of three steps [117]. First, an iterative algorithm is picked/designed for solving an optimization problem. It is preferred that such iterative algorithm is simple, because in the second step, it is unrolled into a neural network. Unrolling stands for representation of computations in one iteration of an algorithm as a layer of a neural network, while the neural network is then build out of several such layers that correspond to the algorithm iterations. Third, a set of neural network parameters to be learned is selected. These may be, for example, step size parameters over the iterations of the corresponding optimization algorithm. An example is given in Section VI-B.
V-D End-to-End Learning
Using deep autoencoders for end-to-end communication is a novel approach with strong potential to improve BER in both single- and multi-antenna systems [12, 118]. Unlike traditional systems that separately design modulation, coding, and decoding to minimize BER and maximize spectral efficiency, autoencoder-based methods treat the entire communication chain as a single neural network. This enables joint optimization from transmitter to receiver by learning an end-to-end mapping that minimizes reconstruction error.
Autoencoders have shown promising performance for various communication scenarios [119, 120, 121, 122, 123, 124]. They have also been applied to radar interference mitigation [125, 126, 127, 128]. In these designs images are used to train the model. Autoencoders have also been applied to target detection [129], focusing on designing and training networks that model background statistics and reconstruct input data while suppressing background-induced peaks. Assuming a known background, this approach emphasizes change detection for dynamic targets, such as people and drones. The dataset used consists of simulated high-range resolution profiles for both background and targets. These profiles represent amplitude variations of the reflected radar signal as a function of range, where peaks correspond to objects or features with significant reflectivity. This approach is conceptually similar to anomaly detection, a common application of autoencoders [35], as the model effectively clusters normal (background) and abnormal (background with target) profiles separately.
In the context of ISAC, several studies are referred to as end-to-end ISAC [130, 131, 132]. However, these approaches typically apply end-to-end learning only to the communication component, while relying on separate methods—often supervised learning—for the radar component. Specifically, in [130], the radar task is detecting the presence of a target and estimating its angle of arrival. The other two papers extend the above approach to the multi-target [131] or multi-user [132] cases. In [133], an attention-based LSTM network is proposed to directly map reflected signal samples to beamforming vectors, capturing temporal correlations to maximize achievable rates. Using an unsupervised learning framework, this approach eliminates intermediate estimation errors, outperforming methods like extended Kalman filtering and deep learning-based two-step designs, while achieving performance close to that of perfect CSI-based beamforming.
In [134], an attention-based LSTM network was proposed for end-to-end predictive beamforming design in ISAC. This scheme captures the temporal correlation in reflected signal samples and directly determines the beamformer, offering a streamlined solution. To address high computational complexity caused by the need to continuously track vehicle motion parameters in vehicular ISAC, the authors in [135] designed an end-to-end predictive beamforming framework based on joint CRLBs. They evaluated sensing performance while ensuring an acceptable downlink communication sum-rate. Similarly, in [136], to adapt precoding policies to dynamic environments, the authors proposed an end-to-end learning method to directly learn downlink multi-user analog and digital hybrid precoders from the received uplink sounding reference signals. Additionally, they developed a parallel proactive optimization network to concurrently learn hybrid precoding.
VI Representative Examples and Performance Evaluation
This section presents two learning-based ISAC waveform and beamformer designs. We describe the problem setups, explain the learning-based methods, and analyze their effectiveness through the following case studies:
  • Case Study I employs unsupervised learning to design an ISAC waveform based on a customized loss function.
  • Case Study II utilizes a modified deep unrolling method to design hybrid analog/digital beamformers for ISAC.
VI-A Case Study I: Unsupervised Learning-Based ISAC Waveform Design
We consider a dual-functional MIMO system, which enables simultaneous downlink communication with multiple users while performing radar sensing. The BS transmits a unified ISAC waveform for both communication and sensing functionalities. The BS communicates with $K$ single-antenna users and detects radar targets concurrently. The BS is also equipped with a uniform linear array (ULA) of $M$ antennas.
VI-A1 The Communication Channel Model
We denote the channel between the BS and the $k$th user by $\mathbf{h}_{k}\in\mathbb{C}^{M}$. We use the Rician distribution to model $\mathbf{h}_{k}$ such that deterministic line-of-sight (LoS) and random non-LoS channel components can be captured as
$\mathbf{h}_{k}$ $=$ $\sqrt{\frac{K_{h_{k}}\eta_{h_{k}}}{K_{h_{k}}+1}}\bar{\mathbf{h}}_{k}+\sqrt{\frac{\eta_{h_{k}}}{K_{h_{k}}+1}}\tilde{\mathbf{h}}_{k}$ (11)
where $K_{h_{k}}$ is the Rician factor, which is the power ratio between the LoS and non-LoS channel components. The large-scale fading coefficient, including the path-loss and log-normal shadowing, is represented by $\eta_{h_{k}}$. In (11), $\bar{\mathbf{h}}_{k}$ captures the LoS channel component, and it can be modeled as
$\bar{\mathbf{h}}_{k}=[1,e^{j2\pi\Delta\sin{\theta_{k}}},\cdots,e^{j2\pi(M-1)\Delta\sin{\theta_{k}}}]$ (12)
where $\theta_{k}$ is the angle-of-departure for user $k$ at the BS, and $\Delta$ is the normalized spacing (by the wavelength) between adjacent antennas. In (11), $\tilde{\mathbf{h}}_{k}$ represents the non-LoS channel components, and its elements are modeled via a complex Gaussian distribution with zero mean and unit variance as $\tilde{\mathbf{h}}_{k}\sim\mathcal{CN}\left(\bm{0}_{{M}\times 1},\mathbf{I}_{M}\right)$. Using (11), the communication channel matrix between the BS antenna array (of $M$ elements) and $K$ single-antenna users can be aggregated as
$\mathbf{H}=[\mathbf{h}_{1},\cdots,\mathbf{h}_{k},\cdots,\mathbf{h}_{K}]^{T}\in\mathbb{C}^{K\times M}.$ (13)
VI-A2 The Sensing Channel
We model the sensing channel between the BS and the target by assuming LoS propagation through the steering vector of the BS antenna array as
$\mathbf{v}(\theta_{t})=[1,e^{j2\pi\Delta\sin{\theta_{t}}},\cdots,e^{j2\pi(M-1)\Delta\sin{\theta_{t}}}],$ (14)
where $\theta_{t}$ is the angle-of-arrival of the target with respect to the receive array of the BS.
VI-A3 ISAC Signal Model
We denote the signal matrix transmitted by the ULA of BS by $\mathbf{X}=[\mathbf{x}_{1},\cdots,\mathbf{x}_{\tau_{d}}]\in\mathbb{C}^{M\times\tau_{d}}$, where $\tau_{d}$ represents the length of communication frame. Hence, the signal matrix $\mathbf{X}$ serves as a dual-functional transmit ISAC waveform. It should be noted that the BS precoder is also embedded in $\mathbf{X}$. We assume that the BS perfectly estimates $\mathbf{H}$ by using uplink pilots. Then, we can write the received signal matrix at the communication users in the downlink as
$\mathbf{Y}=\mathbf{H}\mathbf{X}+\mathbf{N}$ (15)
where $\mathbf{N}=[\mathbf{n}_{1},\cdots,\mathbf{n}_{\tau_{d}}]\in\mathbb{C}^{K\times\tau_{d}}$ is an Gaussian noise matrix having entries modeled as $\mathbf{n}_{i}\sim\mathcal{CN}(\mathbf{0},\sigma^{2}\mathbf{I}_{M}),\forall i$. Next, we denote the desired received signal matrix for $K$ users by $\mathbf{Y}$. Then, we can rewrite (15) as [137]
$\mathbf{Y}=\mathbf{D}+\underbrace{(\mathbf{H}\mathbf{X}-\mathbf{D})}_{\text{MUI}}+\mathbf{N}$ (16)
where $\mathbf{D}$ is the aggregated data matrix for $K$ communication users, and its entries can be drawn from a constellation of any modulation scheme. The second term of (16) represents multi-user interference (MUI). The total power of the MUI term in (16) can be written as
$P_{\text{MUI}}=\|\mathbf{H}\mathbf{X}-\mathbf{D}\|^{2}_{F},$ (17)
where $\|\cdot\|_{F}$ denotes the Frobenius norm.
The optimization of radar beam patterns is fundamentally related to the design of the correlation matrix of the transmit waveform [138, 139]. Hence, the correlation matrix of the ISAC waveform $\mathbf{X}$ can be defined as
$\mathbf{\Sigma}_{X}=\frac{1}{\tau_{d}}\mathbf{X}\mathbf{X}^{H}.$ (18)
Then, the transmit beam pattern of the ISAC waveform $\mathbf{X}$ can be defined through the correlation matrix (18) of the ISAC waveform as [137]
$P(\theta)=\mathbf{v}^{H}(\theta)\mathbf{\Sigma}_{X}\mathbf{v}(\theta),$ (19)
where $\theta$ denotes the detection angle of the target, and $\mathbf{v}(\theta)$ is the array steering vector defined in (14).
VI-A4 ISAC Waveform Optimization
The ISAC waveform $\mathbf{X}$ can be optimized to prioritize communication, sensing, or both functionalities subject to fundamental constraints [137]. Hence, we review the corresponding optimization problem formulations and their optimization-based solutions. Then, we present an alternative ISAC waveform design technique based on unsupervised learning.
  • $\epsilon$-Constraint Method: The optimization problem can be solved in two different ways such that it prioritizes either communication or sensing functionalities.
    First, we consider the case in which communication functionality is prioritized. Then, the $\epsilon$-constraint optimization can be formulated as
    $\underset{\mathbf{X}}{\text{minimize}}$ $\|\mathbf{H}\mathbf{X}-\mathbf{D}\|^{2}_{F}$ (20a)
    subject to $\|\mathbf{X}-\mathbf{X}_{0}\|^{2}_{F}-\epsilon_{\rm sens}\leq 0,$ (20c)
    $\frac{1}{\tau_{d}}\|\mathbf{X}\|^{2}_{F}=P_{T},$
    where $P_{T}$ denotes the ISAC waveform power per frame, and $\mathbf{X}_{0}$ is a reference radar waveform with desirable auto- and cross-correlation properties for target sensing [140, 137]. Specifically, $\mathbf{X}_{0}$ can be designed based on [137, Eq. (9)] for omnidirectional beampatterns and [137, Eq. (15)] for directional ones.
    Alternatively, by prioritizing the sensing functionality, the optimization problem can be formulated as
    $\underset{\mathbf{X}}{\text{minimize}}$ $\|\mathbf{X}-\mathbf{X}_{0}\|^{2}_{F}$ (21a)
    subject to $\|\mathbf{H}\mathbf{X}-\mathbf{D}\|^{2}_{F}-\epsilon_{\rm comm}\leq 0,$ (21c)
    $\frac{1}{\tau_{d}}\|\mathbf{X}\|^{2}_{F}=P_{T}$
    where the goal is to design $\mathbf{X}$ to resemble reference radar waveform $\mathbf{X}_{0}$.
    The above two optimization problems in (20a) and (21a) can be solved optimally using classical convex optimization theory [137, 141, 7].
  • Weighted Sum Method: The waveforms obtained by solving the above two optimization problems may be skewed towards the communications or sensing, respectively. In order to circumvent such circumstances, the ISAC waveforms can be designed by jointly optimizing an objective function computed through a weighted sum of sensing and communication performance metrics by using multi-objective optimization techniques discussed in Section III. This joint design technique focuses on optimizing the fundamental sensing-communication trade-offs. For example, a joint objective can combine the weighted sum of MUI power and the mismatch between the designed ISAC waveform $\mathbf{X}$ and reference waveforms $\mathbf{X}_{0}$. This joint objective function can be then minimized subject to a transmit power constraint as
    $\!\!\!\!\!\!\!\!\!\!\!\!\!\!\!\underset{\mathbf{X}}{\text{minimize}}\!\!\!\!\!\!$ $\eta\|\mathbf{H}\mathbf{X}-\mathbf{D}\|^{2}_{F}+(1-\eta)\|\mathbf{X}-\mathbf{X}_{0}\|^{2}_{F}$ (22a)
    subject to $\frac{1}{\tau_{d}}\|\mathbf{X}\|^{2}_{F}=P_{T},$ (22b)
    where $0\leq\eta\leq 1$ is a weighting factor that controls the trade-off between radar sensing and communication performance.
    The optimization problem (22a)-(22b) can be equivalently reformulated as a quadratically constrained quadratic program. It can also be transformed into a semi-definite program by using the semi-definite relaxation (SDR) techniques, and then solved using standard optimization solvers as in [137].
The computational complexity of such classical convex optimization-based procedures can be prohibitively high when system scales in terms of the numbers of BS antennas, users, and targets. For instance, the computational complexity of the SDR-based optimal solution proposed in [142] for the joint optimization problem in (22a)-(22b) is $\mathcal{O}((M\tau_{d})^{3.5})$.
Efficient learning-based techniques can alternatively be used to provide low-complexity yet sub-optimal ISAC waveform designs with comparable performance [111]. Both supervised and unsupervised learning-based techniques can be applied to solve optimization problems. However, unsupervised learning eliminates the need for labeled data-sets and, thus, is preferable. Specifically, in ISAC systems, collecting labeled data-sets can be both costly and time consuming. Hence, the unsupervised learning-based techniques are more appealing for solveing optimization problems that appear in ISAC waveform designs.
VI-A5 A Novel Unsupervised Learning-based Technique
In learning-based techniques, the design trade-offs between the sensing and communication functionalities can be captured through a custom loss function, which is formulated as a sum of two weighted performance metrics for sensing and communications [111]. For example, minimization of this objective function can be translated into the problem of jointly minimizing the power of MUI for communication users, while retaining the desired correlation properties of the ISAC waveform to detect targets [137]. This joint objective function can be minimized subject to a set of constraints, enforced through multiple Lambda layers, which are custom layers implemented on Pytorch [143, 144, 111].
Lambda layers in DNNs: A Lambda layer in DNNs is a customizable layer. It can be used to embed arbitrary mathematical expressions when constructing sequential functional application programming interfaces (APIs) designed for learning models [143, 144, 111]. The Lambda layers are particularly useful when the intended learning model cannot be implemented by only using standard layers in APIs. They are useful for applying custom feature engineering at the input and enforcing model-specific constraints at the output. Both Pytorch and TensorFlow allow Lambda layers to incorporate arbitrary, user-defined function models into a learning architecture, specifically when standard layers do not provide the required functionality. For instance, in Keras, a Lambda layer can be used to wrap arbitrary expressions or functions as a custom Keras layer, thus enabling execution of specific computations within a custom learning model architecture. In TensorFlow, a Lambda layer can be implemented by using tf.keras.layers.Lambda that wraps a custom function. Pytorch also provides Lambda-like behavior when the nn.Module class is used with a custom forward($\cdot$) method or through inline lambda functions in functional APIs. Lambda layers are useful when implementing and solving constraint optimization problems using deep learning techniques.
A Lambda layer can be used to apply a custom activation function, perform a specific mathematical operation, or implement a custom loss function in designing AI-assisted waveforms for ISAC systems [111]. Through unsupervised learning techniques, an unlabeled data-set can be used to train a DNN by iteratively minimizing the custom loss function subject to constraints introduced by the Lambda layers.
Figure 5: An unsupervised learning-based DNN for ISAC waveforms [111].
TABLE VII: The DNN architecture
Layer Dimension Activation function
Input Layer $2N$
Hidden Layer 1 $20N$ ReLU
Hidden Layer 2 $10N$ ReLU
Hidden Layer 3 $2M\tau_{d}$ Tanh
Output/Lambda Layer $2M\tau_{d}$
Here, we employ an unsupervised learning-based, single-task fully connected DNN to solve the optimization problem (22a)—a spacial form of (4)—and design the waveform. The DNN’s architecture is depicted in Fig. 5, and its structure is furnished in Table VII.
The ISAC waveform design in (22a)-(22b) depends on the desired symbol matrix $\mathbf{D}$, channel matrix $\mathbf{H}$, and the reference waveform $\mathbf{X}_{0}$. Therefore, the input to the DNN consists of $\mathbf{H}$, $\mathbf{D}$, and $\mathbf{X}_{0}$. The input feature vector $\mathcal{F}\in\mathbb{R}^{2N\times 1}$, where $N=K(M+\tau_{d})+M\tau_{d}$, is formed by stacking the real and imaginary parts of the entries of the complex-valued matrices $\mathbf{H}$, $\mathbf{D}$, and $\mathbf{X}_{0}$ in the corresponding long column vectors. The input feature size is thus depends on the number of BS antennas ($M$), number of communication users ($K$), and the communication frame length ($\tau_{d}$).
The output layer consists of a column vector stacked by the real and imaginary parts of the learned ISAC waveform. Specifically, we implement this output layer through a custom Lambda layer to enforce the constraints of the optimization problem (22b) as
$\bm{\vartheta}=\begin{cases}\bm{\vartheta},&\|\bm{\vartheta}\|^{2}_{2}\leq\tau_{d}P_{T}\\ \tau_{d}P_{T}\bm{\vartheta}/\|\bm{\vartheta}\|_{2},&\|\bm{\vartheta}\|^{2}_{2}>\tau_{d}P_{T}.\end{cases}$ (23)
We translate the objective (22a) into the unsupervised learning framework by introducing a custom loss function to handle the sensing-communication trade-off as
$\mathrm{Loss}\!=\!\frac{1}{N_{s}}\!\sum_{n=1}^{N_{s}}\eta\|\mathbf{H}_{n}\mathbf{X}_{n}\!-\!\mathbf{D}_{n}\|_{F}^{2}\!+\!(1\!-\!\eta)\|\mathbf{X}_{n}\!\!-\!{\mathbf{X_{0}}}_{n}\|_{F}^{2},$ (24)
where $N_{s}$ is the mini-batch size. For the $n$th training sample, $\mathbf{H}_{n}$ is the channel matrix, $\mathbf{D}_{n}$ is the desired symbol matrix, ${\mathbf{X_{0}}}_{n}$ is the reference waveform, and $\mathbf{X}_{n}$ is the optimized output waveform.
During the offline training phase, we train the DNN using an unlabeled data-set containing input feature vector samples $\mathcal{F}$. We generated this dateset consisting of $\mathcal{F}$ by following the procedure below. The channel matrices $\mathbf{H}$ are generated according to the channel model in (11) by using Monte-Carlo simulations. The user symbol matrices $\mathbf{D}$ are generated by using a unit-power QPSK modulation scheme. We generate the desired sensing waveform $\mathbf{X}_{0}$ in (22a) and (24) by adhering to ideal omnidirectional/directional waveforms by following [141]. Then, the input feature vector $\mathcal{F}$ can be obtained by stacking the real and imaginary pars of $\mathbf{H}$, $\mathbf{D}$, and $\mathbf{X}_{0}$ for a particular realization. The corresponding dataset can then be generated by stacking multiple samples of input feature vectors $\mathcal{F}$. The mini-batch training is utilized. The DNN model weights can be updated after each batch by applying Adam optimizer. It minimizes the custom loss function (24) via an iterative weight update procedure depicted in Fig. 5 in Pytorch. First, the DNN model is trained and tested by using our dataset. Then, the pre-trained model is used for online prediction of the ISAC waveforms based on new input feature vectors consisting of $\mathbf{H}$, $\mathbf{D}$, and ${\mathbf{X}}_{0}$.
VI-A6 Communication Performance Metrics
We use the downlink achievable sum rate to evaluate the performance of the designed ISAC waveform. The custom cost function aims to minimize the MUI power (17), as the achievable rate is a decreasing function of MUI power. The achievable sum rate can be defined as
$\mathcal{R}_{k}=\sum^{K}_{k=1}\log[2]{1+\gamma_{k}},$ (25)
where $\gamma_{k}$ is the SINR per symbol for the $k$th user in the downlink, and it is given as [137]
$\gamma_{k}=\frac{\mathbb{E}\left[|D_{kq}|^{2}\right]}{\mathbb{E}\left[\left|\mathbf{h}^{T}_{k}\mathbf{x}_{q}-D_{kq}\right|^{2}\right]+\sigma^{2}}.$ (26)
Here $D_{kq}$ is the $(k,q)$th symbol of $\mathbf{D}$, and the expectation is evaluated over symbols [137]. The term $\mathbb{E}[\left|\mathbf{h}^{T}_{k}\mathbf{x}_{q}-D_{kq}\right|^{2}]$ captures the MUI. Since $\mathbb{E}\left[|D_{kq}|^{2}\right]$ is fixed for a given constellation with constant energy, the achievable sum rate can be maximized by minimizing the MUI. Hence, the user SINR (26) and achievable sum rate (25) serve as communication performance metrics of the ISAC waveform designs.
VI-A7 Sensing Performance Metrics
Good correlation properties of ISAC waveforms can enhance sensing success. The correlation matrix of the ISAC waveform is given by (18). The transmit beampattern of the ISAC waveform for a given target direction ($\theta$), computed using (19), can be used to compare ISAC waveforms. The probability of detection, probability of false alarm, and ROC can also be used to evaluate the sensing performance of the designe ISAC waveform. The sensing detector can be designed based on the generalized likelihood ratio test for a binary hypothesis test [145].
VI-A8 Numerical Results and Performance Comparisons
The numerical results and performance comparisons are presented below. The solutions obtained through convex optimization techniques are used as baselines for comparison purposes.
In our example, the carrier frequency is set to 3.2 GHz. The 3GPP urban microcell model [146, Table 5.1] is used to model the path loss, and the Rician factors are set to $K\in[1,3]$. The user-to-BS distance is set to be in $[50,200]$ m range. Unless otherwise specified, $P_{T}=1$ W, $M=\{16,32\}$, $K=4$, and $\tau_{d}=10$ are used. Moreover, a unit power QPSK modulation is adopted to generate the desired symbols in $\mathbf{D}$.
DNN implemented in Pytorch is used. We generate $10^{5}$ samples for the training, testing, and prediction stages. With $N=K(M+\tau_{d})+M\tau_{d}$, for $M=16$ the input layer has $5282N=$ 528 neurons, and the output layer has $M\tau_{d}=160$ neurons. About 60 % and 20 % of the total number samples are used during the offline training and testing phases, respectively. The remaining 20 % of generated samples are used to predict the learning-based ISAC waveforms during the online deployment phase. For training, Adam optimizer with the maximum epoch size of 500 and mini-batch size of 1000 is used. The initial learning rate for the optimizer is set to $0.001$, and the overfitting is circumvented by using early-stopping with patience 20.
Figure 6: Comparison of sum rate between analytical and learning-based ISAC waveform designs for $M=\{16,32\}$, $K=4$, $\tau_{d}=10$, and $\eta=0.2$.
In Fig. 6, a performance comparison between optimization- and learning-based ISAC waveform designs in terms of the achievable sum rate is presented. The sum rate of the ISAC waveform designed by the proposed unsupervised learning-based approach is compared to that of four baseline techniques relying on optimization-based approaches [137]. The channel model from (11) is used, with Rician factors of 1.5, 2.7, 1.2, and 2.5 for the four users. The BS has $M=16$ antennas.
The genie-aided zero-MUI case (Case-I) is included in Fig. 6 to serve as a sum rate upper bound achieved using an ideal ISAC waveform. Case-D and Case-F shown in Fig. 6, we present the sum rate achievable by the proposed unsupervised learning-based technique for omnidirectional and directional desired beampatterns, respectively.
It can be unveiled from Fig. 6 that despite its sub-optimality, the learning-based ISAC waveform provides a comparable sum rate achievable by more complicated optimization-based counterparts. To be specific, by comparing Case-C with Case-E and Case-D with Case-F at a transmit SNR of 10 dB, we observe that the sum rate gaps between the learning-based and analytical baseline waveforms are as small as 0.876 % and 1.314 %, respectively. We also plot the sum rates achievable by the learning-based technique for both omnidirectional and directional ISAC waveforms (Case-G and Case-H) by increasing the number of BS antennas from 16 to 32 to demonstrate the sum rate gains and the scalability of the learning-based ISAC waveform design with respect to the number of BS antennas.
Figure 7: Comparison of target detection ROC performance between learning-based and analytical ISAC waveform designs for $M=\{16,32\}$, $K=4$, $\tau_{d}=10$, and $\eta=\{0.2,0.5\}$.
Figure 7 shows the ROC curves of detection probability ($P_{D}$) versus false alarm probability ($P_{\rm FA}$) for target detection using both optimization- and learning-based ISAC waveform designs. The target detector proposed in [145, Eq. (69)] is used to compare the ROC performance. The ROC curves for Cases-2, 3 and 4 are plotted for the trade-off-based ISAC waveform designs obtained through the unsupervised learning-based approach. As a baseline comparison, in Case-1, the detector ROC curves with an ISAC waveform obtained via the SDR optimization-based solution for (22a)-(22b) in [137] are also plotted. For example, at $P_{\rm FA}=0.5$, the gap between Case-1 and Case-2 in terms of $P_{D}$ is about 0.0119.
Figure 7 also shows how $P_{D}$ can be improved for a given $P_{\rm FA}$ by prioritizing sensing over the achievable rate by suitably varying the weight factor ($\eta$) in the custom loss function (24). For example, by referring to Case-3 and Case-4 at $P_{\rm FA}=0.2$, we observe that $P_{D}$ can be improved by about 7.91 % by decreasing $\eta$ from 0.5 to 0.2. We also show in Fig. 7 that $P_{D}$ can be improved by using a larger array at the BS, compared in Case-2 with Case-3.
In Fig. 8, the transmit sensing beampatterns (19) of the learning-based ISAC waveform are compared with those of the optimization-based baseline techniques in [137]. We consider three target directions for the directional beampattern design as $-\pi/3$, 0, and $\pi/3$. The beampatterns provided by the unsupervised learning-based ISAC waveforms is closely align with the optimal beampatterns derived through SDR optimization-based technique for (22a)-(22b). This observation concludes that low-complexity learning-based approaches can be beneficial in circumventing the high computational complexity of optimization-based ISAC waveform designs.
VI-A9 Computational Complexity and Design Insights
The computational complexity of learning- and optimization-based ISAC waveforms can be compared in terms of floating-point operations. For the optimization-based ISAC waveform design with a sensing-centric omnidirectional beampattern, the complexity scales as $\mathcal{O}(MK\tau_{d}+M\tau_{d}^{2})$, due to two matrix multiplications and one SVD operation [137]. The complexity of the optimization-based ISAC waveform design with sensing-centric directional beampattern scales with $\mathcal{O}(M{\tau_{d}}^{2}+M^{2}\tau_{d}+MK\tau_{d}+M^{3}+M^{2}K)$ as it has one Cholesky decomposition, one SVD operation, and four matrix multiplications. The computational complexity of the trade-off based joint ISAC waveform design in (22a)-(22b) scales as $\mathcal{O}((M\tau_{d})^{3.5})$ when it is solved by leveraging SDR optimization-based technique in [142].
The computational complexity of a fully-connected layer scales as $\mathcal{O}((2D_{I}-1)D_{O})$, where $D_{I}$ and $D_{O}$ are the input and output dimensions, respectively [147]. In the worst-case scenario, they depend on the dimension of the input space $\mathcal{F}\in\mathbb{R}^{2N\times 1}$, where $N=K(M+\tau_{d})+M\tau_{d}$. During the ISAC waveform prediction stage, the overall computational complexity of the unsupervised learning-based technique scales as $\mathcal{O}(KM^{2}\tau_{d}+KM\tau^{2}_{d}+M^{2}\tau^{2}_{d})$.
Figure 8: Comparison of beampatterns between analytical and learning-based ISAC waveform designs for $M=32$, $K=4$, $\tau_{d}=10$, and $\eta=0.2$.
VI-B Case Study II: Hybrid Beamforming Using Learned
Optimizer
We start by presenting the alternating optimization-based method from [148] for designing hybrid beamforming (HBF), which incorporates a modified deep unrolling method. Specifically, it leverages a detection network (DetNet) [149, 150] to unroll the iterations of the projected gradient descent algorithm into a neural network. In addition, we provide a numerical study that evaluates the spectral efficiency performance of the HBF method.
VI-B1 Signal Model
Consider a point-to-point communication system in which the BS and the mobile station are equipped with $N_{t}$ and $N_{r}$ antennas. The post-processed signal at the BS can be expressed as [148]
$\mathbf{y}=\sqrt{\rho}\mathbf{V}^{H}\mathbf{H}\mathbf{F}_{{\mathrm{RF}}}\mathbf{F}_{{\mathrm{BB}}}\mathbf{s}+\mathbf{V}^{H}\mathbf{n},$ (27)
where $\rho$ is the average received power, $\mathbf{V}\in\mathbb{C}^{N_{r}\times N_{s}}$ is the digital combiner, $\mathbf{F}_{\mathrm{RF}}\in\mathbb{C}^{N_{t}\times N_{\mathrm{RF}}}$ is analog beamformer, $\mathbf{F}_{\mathrm{BB}}\in\mathbb{C}^{{N_{\mathrm{RF}}}\times N_{s}}$ is digital baseband precoder, $\mathbf{H}$ is channel matrix, and $\mathbf{n}$ is an AWGN vector with zero mean and variance $\sigma^{2}_{n}$. Here $N_{\mathrm{RF}}$ is the number of RF chains in the BS, $N_{s}$ is the number of data streams, and $\mathbf{s}$ is the signal vector with $\mathbb{E}[\mathbf{s}\mathbf{s}^{H}]=\mathbf{I}$. The achievable spectral efficiency for Gaussian symbols is given by [151]
$\!\!\!\!R=\mathrm{log}_{2}\det\left(\mathbf{I}_{{N_{\mathrm{s}}}}+\frac{\rho}{\sigma_{\mathrm{n}}^{2}N_{\mathrm{s}}}\mathbf{V}^{\dagger}\mathbf{H}\mathbf{F}_{\mathrm{RF}}\mathbf{F}_{\mathrm{BB}}\mathbf{F}_{\mathrm{BB}}^{H}\mathbf{F}_{\mathrm{RF}}^{H}\mathbf{H}^{H}\mathbf{V}\right),$ (28)
where $(\cdot)^{\dagger}$ is the matrix pseudo-inverse. The spectral efficiency maximization can be approximately achieved by [151]
[…]
where $f_{mn}^{\mathrm{RF}}$ is the $(m,n)$th element of $\mathbf{F}_{\mathrm{RF}}$, and $\mathbf{F}_{\mathrm{opt}}$ is the unconstrained optimal digital precoder.
VI-B2 Hybrid Beamforming Design
We now introduce the alternating optimization-based method proposed in [148] for HBF, along with the concept of rolling the projected gradient descent algorithm, inspired by DetNet, which we test here.
The main idea of alternating optimization is to optimize $\mathbf{F}_{\mathrm{RF}}$ with $\mathbf{F}_{\mathrm{BB}}$ given and the second constraint in (29) omitted. Then, $\mathbf{F}_{\mathrm{BB}}$ is optimized to meet the constraint given the optimized $\mathbf{F}_{\mathrm{RF}}$. Thus, the following problem is first considered:
[…]
The vectorization operator is performed on the matrices $\mathbf{F}_{\mathrm{opt}}$, $\mathbf{F}_{\mathrm{RF}}$, and $\mathbf{F}_{\mathrm{BB}}$, resulting in $\tilde{\mathbf{z}}=\text{vec}(\mathbf{F}_{\mathrm{opt}})$, $\tilde{\mathbf{x}}=\text{vec}(\mathbf{F}_{\mathrm{RF}})$, and $\tilde{\mathbf{B}}=\mathbf{F}_{\mathrm{BB}}\otimes\mathbf{I}_{N_{t}}$, where $\otimes$ denotes the Kronecker product. Then the resulting complex-valued quantities are transformed into real-valued ones, denoted as
$\mathbf{z}\triangleq\begin{bmatrix}\mathfrak{R}(\tilde{\mathbf{z}})\\ \mathfrak{I}(\tilde{\mathbf{z}})\end{bmatrix},\mathbf{x}\triangleq\begin{bmatrix}\mathfrak{R}(\tilde{\mathbf{x}})\\ \mathfrak{I}(\tilde{\mathbf{x}})\end{bmatrix},\mathbf{B}\triangleq\begin{bmatrix}\mathfrak{R}(\tilde{\mathbf{B}})&-\mathfrak{I}(\tilde{\mathbf{B}})\\ \mathfrak{I}(\tilde{\mathbf{B}})&\mathfrak{R}(\tilde{\mathbf{B}})\end{bmatrix},$ (31)
where $\mathfrak{R}(\cdot)$ and $\mathfrak{I}(\cdot)$ denote the the real and imaginary parts of a complex-valued vector (matrix), respectively.
Thus, problem (30) can be reformulated as a least-squared problem, denoted as
$\mathbf{x}_{\mathrm{ML}}=\underset{{\mathbf{x}\in\mathcal{S}}}{\operatorname*{\operatorname*{\operatorname*{argmin}}}}\left\|\mathbf{z}-\mathbf{B}\mathbf{x}\right\|^{2},$ (32)
where $\mathcal{S}$ is defined as
$\mathcal{S}\triangleq\{\mathbf{x}:x_{i}+jx_{N_{\mathrm{t}}N_{\mathrm{RF}}+i}=\tilde{x}_{i},|\tilde{x}_{i}|=1,i=1,\ldots,N_{\mathrm{t}}N_{\mathrm{RF}}\}.$
This constrained least-squares problem can be solved using DetNet. Specifically, the output of $l$th layer of the neural network ($\mathbf{x}_{l}$) can be expressed as
[…]
where $\delta_{\ell}$ is the stepsize and $\Pi_{\ell}$ nonlinear projection operator, which are learned by proposed DNN architecture (see architecture details in [148]). The employed loss function is
$\mathcal{L}=\sum_{\ell=1}^{L}\log{\ell}\left\|\mathbf{z}-\mathbf{B}\mathbf{x}_{\ell}\right\|^{2},$ (34)
where $L$ is the number of layers of the unrolled network.
Once offline training is completed, the well-trained neural network is used to perform online HBF design. Then, $\mathbf{F}_{{\mathrm{RF}}}$ can be reconstructed from $\mathbf{x}_{L}$. Define $\tilde{\mathbf{H}}=\mathbf{H}\mathbf{F}_{{\mathrm{RF}}}$ and $\mathbf{Q}=\mathbf{F}^{H}_{{\mathrm{RF}}}\mathbf{F}_{{\mathrm{RF}}}$. The digital precoder design problem can be expressed as [148]
[…]
which has the well-known water-filling solution:
$\mathbf{F}_{{\mathrm{BB}}}=\mathbf{Q}^{{-\frac{1}{2}}}\tilde{\mathbf{U}}\tilde{\bm{\Gamma}},$ (36)
where the columns of $\tilde{\mathbf{U}}$ are taken from the right singular vectors corresponding to the $N_{s}$ largest singular values of $\tilde{\mathbf{H}}\mathbf{Q}^{{-\frac{1}{2}}}$, and $\tilde{\bm{\Gamma}}$ is a diagonal matrix whose elements are de- fined by the power allocated to the $N_{s}$ data streams.
VI-B3 Numerical Results
In this section, the performance of the rolled neural network tailored to the HBF design is tested. The system parameters are set as $N_{t}=32$, $N_{s}=N_{\mathrm{RF}}=2$. The channel $\mathbf{H}$ is modeled as Rayleigh fading channel. The sizes of datasets for training and testing are 500 and 100. For training phase, Adam optimizer with a learning rate of 0.01 is used. In Figs. 10 and 10, the HBF design using unrolled neural network is named as ‘unrolled HBF’, and unconstrained fully DBF as ‘optimal DBF’.
The spectral efficiency versus the SNR for $N_{r}=16$ and $N_{r}=32$ are shown in Fig. 10. It is observed that for both cases, the spectral efficiency gaps between optimal DBF and unrolled HBF for $\text{SNR}=-10\,\text{dB}$ is smaller than that for $\text{SNR}=15\,\text{dB}$. In addition, the spectral efficiency gap between optimal DBF and unrolled HBF for $N_{r}=32$ is slightly smaller than that for $N_{r}=16$ across the whole SNR region. This implies that the spectral efficiency performance of unrolled HBF scheme is close to that of optimal BF when the number of receiving antennas is large and the SNR is in low region.
Next, the spectral efficiency performance versus the number of receiving antennas is tested in the setting where the SNR is $10$ dB. As shown in Fig. 10, as the number of receiving antennas increases, the spectral efficiency gap between optimal BF and unrolled HBF gets smaller, which means that the unrolled HBF scheme in our context performs well when the number of receiving antennas is large.
1010-10- 1088-8- 866-6- 644-4- 422-2- 20022224444666688881010101012121212141414141616161655551010101015151515SNR (dB)Spectral Efficiency (bits/s/Hz)Optimal DBF, $N_{r}=32$Unrolled HBF, $N_{r}=32$Optimal DBF, $N_{r}=16$Unrolled HBF, $N_{r}=16$
Figure 9: Spectral efficiency versus SNR for $N_{t}=32$, $N_{s}=N_{\mathrm{RF}}=2$, $N_{r}=16$ and $N_{r}=32$.
[…]
Figure 10: Spectral efficiency versus the number of receiving antennas for $N_{t}=32$, $N_{s}=N_{\mathrm{RF}}=2$, SNR is $10$ dB.
VII Future Research Directions
The range of problems in ISAC, which have been shown promise to benefit from AI is wide. Here, we aim to identify several promising directions for future research in the area of AI-empowered ISAC. This list is not intended to be complete, but it focuses more on ISAC open problems, particularly in waveform and beamforming design.
VII-A End-to-End ISAC
While AI-based constellation design for interference mitigation in communications has been under consideration recently [126, 123, 13], there is currently no work for ISAC. This is due to the unique challenges posed by ISAC, where a unified waveform is used for both communication and sensing tasks, leading to overlapping signal components in temporal, frequency, and spatial domains. Future directions in end-to-end learning-based ISAC can focus on addressing this problem. In addition, addressing practical challenges, such as minimizing symbol error rate in the presences of interference in various forms, including self-interference, inter-cell interference, and inter-user interference. Further, extending these methods to multi-user and multi-target scenarios, while enhancing spectral efficiency and ensuring scalability, will be critical to realizing the full potential of ISAC systems in practical deployments.
While current approaches [130, 131, 132] have demonstrated promising path in jointly optimizing sensing and communication tasks, their performance, particularly in realistic scenarios with interference, remains unknown. These works also ignore interference in its various forms including self-interference [152, 153, 154], inter-cell interference[155], and inter-user interference [156], among others. Future research should explore advanced techniques to integrate interference mitigation into end-to-end learning frameworks for both single-carrier and multi-carrier methods, including OFDM and OTFS. Problem of interference reduction involves using end-to-end learning to develop new, flexible constellations that are specifically designed to be aware of self-interference. To achieve this, the research will leverage autoencoder-based techniques for self-interference cancellation, which is a promising approach since interference can be identified at the transmitter. To address this challenge, this task will utilize unsupervised learning techniques to develop adaptive constellations for ISAC.
VII-B Transformer Neural Networks for ISAC Design
Intelligent vehicular networks are a key use case of ISAC in 6G [157]. Classical waveform and beamformer designs, such as beam codebooks and analytical optimizations, may fall short in high-speed vehicular environments. To address this, AI-driven ISAC designs can leverage multimodal sensory data—such as radar, vision, and LiDAR—to enhance waveform and beamforming performance under challenging propagation conditions [30]. For example, the high-resolution images from cameras connected to roadside BSs, point clouds from LiDAR, reflected electromagnetic signals from transmitted waves, localization information from GPS-enabled sensors can be collected and used to enhance the design of ISAC waveforms. In [158], a transformer-based beamforming framework is proposed that leverages 3D ResNet-18 to extract and fuse multi-modal features (vision, radar, LiDAR, and position) for sensing-assisted communication in dynamic and complex road conditions.
AI techniques can extract 3D features from multimodal inputs such as vision, radar, LiDAR, and position data. Transformer networks [159], the foundation of LLMs like ChatGPT and Gemini, can fuse these features to learn ISAC waveforms and beamformers. Their masked multi-head attention preserves structural dependencies, while customized masking [160] enables embedding of sensing and communication-specific side information. These capabilities make transformers well-suited for addressing complex ISAC optimization problems with improved accuracy and reduced data requirements.
VII-C Digital Twin-aided ISAC System Designs
Digital twins are expected to play a key role in enabling 6G wireless technologies, offering proactive analytics and efficient training of AI models to replicate real-world propagation environments [161]. A digital twin comprises pre-trained models that support real-time decision-making for multi-functional wireless services.
In the context of ISAC, digital-twin-aided architectures are particularly promising for dynamic and distributed scenarios. By collecting spatial and temporal information from distributed access points and sensors, digital twins can construct 3D models of the surrounding environment, capturing positions, orientations, and shapes of objects such as scatterers, targets, and users. These digital replicas enable simulation of communication and sensing channels, which can be used to design ISAC transmission strategies and train AI models for unified waveform and beamformer design.
Advanced AI tools in data analysis, pattern recognition, and real-time reasoning further enhance digital twin capabilities for ISAC tasks—ranging from waveform and beamformer optimization to spectral-efficient constellations, beam pattern generation, interference mitigation, and resource allocation. A key challenge is making digital twins self-configurable with online learning capabilities. However, realizing practical digital twins for ISAC demands substantial engineering effort to accurately replicate complex physical environments using fast and efficient AI models.
VIII Conclusions
This tutorial provides a comprehensive overview of AI-driven ISAC systems, emphasizing their significance in 6G networks. It explores fundamental ISAC concepts, traditional and learning-based optimization techniques, and AI-enhanced waveform and beamforming design. Through case studies, it demonstrates AI’s potential in optimizing ISAC performance, highlighting trade-offs and computational efficiency. Lastly, it outlines future research directions, bridging traditional signal processing with modern AI advancements to guide researchers and practitioners in advancing ISAC technology.